### v1

# Thinking Outside the (Gray) Box: A Context-Based Score for Assessing Value and Originality in Neural Text Generation

Abstract
Despite the increasing use of large language models for creative tasks, their outputs often lack diversity. Common solutions, such as sampling at higher temperatures, can compromise the quality of the results. Drawing on information theory, we propose a context-based score to quantitatively evaluate value and originality. This score incentivizes accuracy and adherence to the request while fostering divergence from the learned distribution. We propose using our score as a reward in a reinforcement learning framework to fine-tune large language models for maximum performance. We validate our strategy through experiments in poetry generation and math problem solving, demonstrating that it enhances the value and originality of the generated solutions.
Thinking Outside the (Gray) Box: A Context-Based Score for
Assessing Value and Originality in Neural Text Generation
Giorgio Franceschelli University of Bologna giorgio.franceschelli@unibo.it                        Mirco Musolesi University College London University of Bologna m.musolesi@ucl.ac.uk
1 Introduction
Foundation models Bommasani et al. (2021), particularly large language models (LLMs) Bubeck et al. (2023); Gemini Team et al. (2023); Touvron et al. (2023), are significantly transforming creative activities. They can serve as a foundation for co-creation systems involving human and artificial authors Lin et al. (2023); they can be utilized to generate software code Rozière et al. (2023); they can even be employed to foster scientific research Boiko et al. (2023). However, the nature of the self-supervised learning algorithms used for the training of these models tend to make them to generate samples as close as possible to the training data distribution Franceschelli and Musolesi (2024a). In addition, fine-tuning, such as that based on reinforcement learning from human feedback (RLHF) Christiano et al. (2017), is often necessary to generate appropriate and accurate responses. However, this process tends to further reduce output diversity Kirk et al. (2024), and linguistic creativity tends to be lower than that of humans Lu et al. (2025). On the contrary, LLMs for creative tasks should produce more novel and surprising texts that maintain a high level of correctness and adherence to the request. One typical solution is to sample at higher temperature to increase diversity. However, this might lead to generating incoherent text Peeperkorn et al. (2024).
In order to address the issues described above, we propose a new training approach for creative tasks based on CoVO, a Context-based score for Value and Originality, with the goal of taking into consideration both value and originality of the neurally-generated text in the optimization of LLMs. The definition of CoVO is grounded in the analysis of mutual information MacKay (2003) between the model’s outputs and inputs, and vice versa. More specifically, we formulate a new optimization problem where, given a specific input, the desired output is derived by simultaneously maximizing the conditional probability of the input given the output and minimizing the conditional probability of the output given the input under the generative model. In this way, we optimize for solutions that are appropriate for the input request but also different from the outputs we would normally obtain from the model. In particular, we show that our information-theoretic score can be used as a reward in RL-based fine-tuning algorithms, guiding pre-trained models toward more diverse yet valuable solutions.
We also present the theoretical foundations of our approach and we discuss several strategies for the practical implementation of the proposed approach and its adaptation to current LLMs. Experiments on math problem solving and poetry generation show that our score correlates with domain-specific measures for value and originality, and that our approach can increase the quality and diversity of outputs, presenting itself as a candidate for creativity applications of current foundation models.
The remainder of the paper is structured as follows. First, we provide an overview of the state of the art in this area (Section 2) and introduce the key concepts and background of our work (Section 3). Then, we discuss the definition of CoVO (Section 4) and show different ways to use it in practice as a reward for fine-tuning autoregressive models based on RL (Section 5). We validate the proposed solutions on poetry generation and math problem solving (Section 6). Finally, we discuss future work and limitations of our approach (Sections 7 and 8).
2 Related Work
2.1 Information Theory and Creativity
The quest to provide a mathematical and computational definition of creativity has been a significant focus in recent decades. Numerous methods have been developed to define various dimensions or attributes for evaluating the creativity of AI-generated products (see, for example, Franceschelli and Musolesi, 2024a). However, these methods are often domain-specific and typically require substantial human effort to implement and assess. In contrast, solutions based on information theory Shannon (1948); Cover (1999) offer a more universally applicable approach.
Information-theoretic methods can quantify creativity by measuring the novelty and complexity of generated outputs, without the need for extensive human intervention, making them suitable for a wide range of domains. Bayesian surprise Baldi and Itti (2010), i.e., the divergence between a prior and a posterior belief, has been extensively used to measure either novelty França et al. (2016); Varshney et al. (2019) or surprise Mazzaglia et al. (2022); Schmidhuber (2010). Nevertheless, Varshney (2019) demonstrated that there is a mathematical limit for Bayesian surprise when combined with quality measures. Surprisal Tribus (1961), i.e., Shannon’s self-information, has also been used Bunescu and Uduehi (2019); Fernandez Monsalve et al. (2012); Barto et al. (2013) extensively discuss surprisal and Bayesian surprise, and how novelty differ from them. Crucially, in the context of RL, surprisal has been used as a form of intrinsic motivation to encourage the agent to explore more Achiam and Sastry (2017). Sun et al. (2025) apply this idea to improve exploration in RLHF Christiano et al. (2017). Burns (2006) proposes to use entropy for expectation and violation, plus posterior probability for explanation in the context of aesthetic experience. Additionally, mutual information has been applied to neural conversation models to improve both diversity and appropriateness Li et al. (2016). However, all these existing approaches are not able to capture and simultaneously optimize value and originality at the same time.
2.2 LLMs and Creativity
Since the introduction of GPT models Brown et al. (2020); OpenAI (2023) and their competitors (e.g., Touvron et al., 2023), researchers have been keenly exploring the potential for LLMs to exhibit creativity and the methods to achieve this Franceschelli and Musolesi (2024b). For example, human creativity tests like the Alternate Uses Test have been employed to evaluate the creativity of LLMs Stevenson et al. (2022) and to investigate methods for enhancing their performance Goes et al. (2023); Summers-Stay et al. (2023). Porter and Machery (2024) report that non-expert poetry readers already favor AI-generated poems over human-authored ones. In contrast, Davis (2024) argues that ChatGPT’s poetry is incompetent and banal. Either way, instead of being used off-the-shelf, LLMs can be fine-tuned to produce more rhyming poems Popescu-Belis et al. (2023) or utilized in zero-shot settings to emulate the writing styles of famous authors Sawicki et al. (2023). It has also been shown that these models can be fine-tuned via RLHF Christiano et al. (2017) to write short poems that human evaluators find more creative Pardinas et al. (2023). Finally, it is possible to leverage quality-diversity algorithms to generate more creative products; these methods can be based on human Ding et al. (2023) or AI Bradley et al. (2024) feedback to measure the quality of the generated outputs.
3 Preliminaries
3.1 Language Modeling
A $\boldsymbol{\theta}$-parameterized autoregressive language model is a probability distribution $p_{\boldsymbol{\theta}}(\mathbf{x})$ over a variable-length text sequence $\mathbf{x}=(x_{1}\ldots x_{T})$, where $T$ is the sequence length and each token $x_{t}$ is in a finite vocabulary $\mathcal{V}$ of size $N$. The probability distribution is factorized as $p_{\boldsymbol{\theta}}(\mathbf{x})=\prod_{t=1}^{T}p_{\boldsymbol{\theta}}(x_{t}|\mathbf{x_{<t}})$, where $\mathbf{x_{<t}}=x_{1}\ldots x_{t-1}$. The language model is usually trained to maximize the likelihood of the true distribution $p^{*}(\mathbf{x})$ for any $\mathbf{x}$ from a reference dataset (the training set). In other words, given an input $\mathbf{x_{<t}}$, the model learns to approximate the probability of each token from $\mathcal{V}$ being $x_{t}$. While this makes such a model immediately capable of scoring the probability of a given text, it also allows for the generation of new sentences. Given a conditional input (prompt) $\mathbf{z}=(z_{1}\ldots z_{L})$, we can decode $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})$ as the continuation of $\mathbf{z}$, i.e., through the factorized representation $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})=\prod_{t=1}^{T}p_{\boldsymbol{\theta}}(x_{t}|\mathbf{x_{<t}},\mathbf{z})$.
3.2 Reinforcement Learning for Language Models
Due to its adherence to the formal framework of Markov decision processes Sutton and Barto (2018), RL can be used as a solution to the generative modeling problem in the case of autoregressive tasks such as text generation Bachman and Precup (2015). The LLM plays the role of the agent, and each generated token represents an action $a_{t}$. The current version of the generated output $\mathbf{x_{t}}$ is part of the state $\mathbf{s_{t}}$ (potentially with additional information such as initial prompts). Finally, the reward $r_{t+1}$ measures the “quality” of the current output. A common strategy is to assign a zero reward for each $\mathbf{x}_{t},t\neq T$ and a sentence-based reward when the final output is generated. Within this framework, any policy-based method, such as Proximal Policy Optimization (PPO) Schulman et al. (2017), can be employed to train or fine-tune the LLM to optimize a given objective. Indeed, RL facilitates the use of non-differentiable reward functions, enabling the optimization of test-time metrics, domain-specific targets, and human preferences Franceschelli and Musolesi (2024c). To avoid deviating too much from the pre-trained model while also encouraging exploration, the full reward function usually considers a Kullback-Leibler (KL) divergence term as follows Stiennon et al. (2020):
$R(\mathbf{z},\mathbf{x})=r(\mathbf{z},\mathbf{x})-\beta\log\frac{p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})}{p_{\boldsymbol{\theta}^{PT}}(\mathbf{x}|\mathbf{z})},$ (1)
where $\boldsymbol{\theta}^{PT}$ is the set of parameters of the pre-trained model, $\beta$ weighs the KL penalty, and $r(\mathbf{z},\mathbf{x})$ can be any non-differentiable function to score the model’s output.
In particular, RLHF Christiano et al. (2017) has been widely adopted as the final training stage in various popular LLMs to align them with human preferences, i.e., by using a learned reward function that approximates human feedback. While highly effective, RLHF suffers from several open problems Casper et al. (2023). These include obtaining accurate, unbiased, and representative human feedback, as well as making the process resource-efficient. Moreover, RLHF tends to be complex and unstable due to the additional requirement of training a reward model and its reliance on RL. To address these challenges, Direct Preference Optimization (DPO) Rafailov et al. (2023) directly optimizes the language model without the need for explicit reward modeling or RL. DPO implicitly performs reward maximization with a KL-divergence penalty through the following loss:
$\mathcal{L}_{DPO}=\min(-\mathbb{E}_{(\mathbf{z},\mathbf{x_{w}},\mathbf{x_{l}})\sim X}[\log\sigma($ $\beta\log\frac{p_{\boldsymbol{\theta}}(\mathbf{x_{w}}|\mathbf{z})}{p_{\boldsymbol{\theta}^{PT}}(\mathbf{x_{w}}|\mathbf{z})}$ (2)
-- $\beta\log\frac{p_{\boldsymbol{\theta}}(\mathbf{x_{l}}|\mathbf{z})}{p_{\boldsymbol{\theta}^{PT}}(\mathbf{x_{l}}|\mathbf{z})})])$
where $\sigma$ is the sigmoid function, $\beta$ is a parameter controlling the deviation from the base model, and $\mathbf{x_{w}},\mathbf{x_{l}}$ are the chosen and rejected outputs, respectively. In other words, DPO fits an implicit reward whose optimal policy is the language model itself.
4 A Context-Based Score for Valuable and Original Generation
Our goal is to derive a score that is able to quantify both value and originality at the same time. As discussed in depth by Csikszentmihalyi (2014), creativity depends on the context in which the product is created, as the context provides the task identification and the domain information necessary to generate and validate the outcome. In turn, the output aims to solve the given task and provide a meaningful, original contribution to the current domain. Thus, our proposed score has its roots in mutual information, which represents a quantitative way to study the relationship between contextual, prior information and a produced posterior outcome. More specifically, we start from the (point-wise) mutual information between two variables $x$ and $y$:
$I(x,y)=h(x)-h(x|y)=h(y)-h(y|x)$ (3)
where the entropy is $h(a)=-\log p(a)$, therefore:
$I(x,y)$ $=\log p(x|y)-\log p(x)$ (4)
$=\log p(y|x)-\log p(y).$
Let us now assume $x$ to be our input vector $\mathbf{x}$ and $y$ our output vector $\mathbf{y}$, obtaining:
$I(\mathbf{x},\mathbf{y})=\log p(\mathbf{y}|\mathbf{x})-\log p(\mathbf{y}).$ (5)
We can generalize $I(\mathbf{x},\mathbf{y})$ with two scaling factors:
$I(\mathbf{x},\mathbf{y},\lambda_{1},\lambda_{2})=\lambda_{1}\log p(\mathbf{y}|\mathbf{x})-\lambda_{2}\log p(\mathbf{y}),$ (6)
where $I(\mathbf{x},\mathbf{y})$ is just $I(\mathbf{x},\mathbf{y},1,1)$. By applying the Bayes theorem, i.e., $\log p(a|b)=\log p(b|a)+\log p(a)-\log p(b)$, we can substitute the $\log p(\mathbf{y})$ term as follows:
$I(\mathbf{x},\mathbf{y},\lambda_{1},\lambda_{2})$ $=$ $\lambda_{1}\log p(\mathbf{y}|\mathbf{x})-\lambda_{2}\log p(\mathbf{y}|\mathbf{x})-$ (7)
$\lambda_{2}\log p(\mathbf{x})+\lambda_{2}\log p(\mathbf{x}|\mathbf{y})$
$=$ $(\lambda_{1}-\lambda_{2})\log p(\mathbf{y}|\mathbf{x})+$
$\lambda_{2}\log p(\mathbf{x}|\mathbf{y})-\lambda_{2}\log p(\mathbf{x}).$
Since our goal is to find the optimal $\mathbf{y}$ for a given $\mathbf{x}$, the last term can be ignored. Moreover, we now define $\lambda_{v}=\lambda_{2}$ and $\lambda_{o}=\lambda_{2}-\lambda_{1}$, thus obtaining the following objective:
$\overline{\mathbf{y}}=\underset{\mathbf{y}}{\text{argmax}}(\lambda_{v}\log p(\mathbf{x}|\mathbf{y})-\lambda_{o}\log p(\mathbf{y}|\mathbf{x})).$ (8)
Let us now consider the case where $\lambda_{v},\lambda_{o}>0$, for example, $\lambda_{v}=\lambda_{o}=1$. Solving this maximization problem involves finding the target $\mathbf{y}$ that maximizes the posterior probability of $\mathbf{x}$ while also being unlikely given $\mathbf{x}$. In other words, the optimal $\mathbf{y}$ must be unexpected and diverse from $p(\mathbf{y}|\mathbf{x})$, but it must also be explainable by $\mathbf{x}$. While $-\log p(\mathbf{y}|\mathbf{x})$, commonly known as surprisal Tribus (1961), is widely used to measure diversity and surprise Barto et al. (2013), the other term, $\log p(\mathbf{x}|\mathbf{y})$, can be used to measure value and effectiveness. Indeed, if the request (e.g., a problem or a task) can be easily predicted from the outcome, the outcome must be a (good) example of that task or a correct solution for that problem. While other prominent definitions such as Boden’s Boden (2003) seek to include a third requirement for creativity, i.e., novelty, we believe that by considering only a specific context, novelty and surprise become indistinguishable. In particular, being novel in a specific context $\mathbf{x}$ means doing something different from what has been previously experienced under $\mathbf{x}$ by the creator. However, since a self-supervised-trained model can be seen as a compression of training data Franceschelli et al. (2024), $-\log P(\mathbf{y}|\mathbf{x})$ becomes not only a measure of unexpectedness but also of novelty. Therefore, we choose to refer to this term as originality, as it encompasses both novelty and surprise. On the other hand, $\log P(S|T)$ is less ambiguous and maps directly into value.
We now provide a formal definition of our proposed score. In summary, the CoVO (Context-based Value and Originality) score for a target $\mathbf{y}$ given a source $\mathbf{x}$ on a reference probability distribution $p$ is defined as:
CoVO Score $s_{CoVO}(\mathbf{x},\mathbf{y},p)=\underbrace{\lambda_{v}\log p(\mathbf{x}|\mathbf{y})}_{\textrm{Value}}\underbrace{-\lambda_{o}\log p(\mathbf{y}|\mathbf{x})}_{\textrm{Originality}}$ (9)
5 Implementation and Optimization with Autoregressive Models
We now discuss the implementation of the CoVO score with autoregressive models. Using the notation introduced in Section 3.1, in the context of a $\boldsymbol{\theta}$-parameterized LLM, $p(\mathbf{y}|\mathbf{x})$ can be expressed as $\prod_{t=1}^{T}p_{\boldsymbol{\theta}}(y_{t}|\mathbf{y}_{<t},\mathbf{x})$. However, considering just the product of all the conditioned probabilities for an optimization problem would lead to preferring shorter sequences. To avoid this, we propose to use the $T$-th root: $T]{\prod_{t=1}^{T}p_{\boldsymbol{\theta}}(y_{t}|\mathbf{y}_{<t},\mathbf{x})}$. By leveraging the properties of the logarithm, we obtain:
$s_{CoVO}^{AR}$ $=\lambda_{v}\frac{\sum_{i=1}^{|\mathbf{x}|}\log p_{\boldsymbol{\theta}}(x_{i}|\mathbf{x}_{<i},\mathbf{y})}{|\mathbf{x}|}$ (10)
$-\lambda_{o}\frac{\sum_{j=1}^{|\mathbf{y}|}\log p_{\boldsymbol{\theta}}(y_{j}|\mathbf{y}_{<j},\mathbf{x})}{|\mathbf{y}|}.$
It is worth noting that the vocabulary of an LLM can be extremely large, which can cause $p_{\boldsymbol{\theta}}(a|b)$ to be small even when $a$ is the most probable event given $b$. In particular, when an LLM generates $\mathbf{y}$ given $\mathbf{x}$ and then evaluates both $p_{\boldsymbol{\theta}}(\mathbf{y}|\mathbf{x})$ and $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{y})$, this can lead to a significant discrepancy between the magnitude of value and diversity. Since $\mathbf{y}$ has been sampled from $p_{\boldsymbol{\theta}}$, its probability would be high by definition. However, there may be various ways (possibly through synonyms) to define $\mathbf{y}$, leading to a smaller probability of $\mathbf{x}$.
Inspired by Macedo et al. (2004), we propose to normalize $p_{\boldsymbol{\theta}}(a|b)$ via $n^{\prime}=\frac{n-n_{min}}{n_{max}-n_{min}}$. For probabilities, $n_{min}=0$, while $n_{max}=\max p_{\boldsymbol{\theta}}(b)$, thus obtaining the overall mapping for $p_{\boldsymbol{\theta}}$: $\frac{p_{\boldsymbol{\theta}}(y_{t}|\mathbf{y}_{<t},\mathbf{x})}{\max p_{\boldsymbol{\theta}}(\mathbf{y}_{<t},\mathbf{x})}$. Once again, by applying the properties of logarithms, we obtain:
CoVO Score with Autoregressive Models $s_{CoVO}^{AR_{norm}}(\mathbf{x},\mathbf{y},p_{\boldsymbol{\theta}})=\lambda_{v}s_{v}(\mathbf{x},\mathbf{y},p_{\boldsymbol{\theta}})+\lambda_{o}s_{o}(\mathbf{x},\mathbf{y},p_{\boldsymbol{\theta}})=$ (11) $\lambda_{v}\frac{\sum_{i=1}^{|\mathbf{x}|}(\log p_{\boldsymbol{\theta}}(x_{i}|\mathbf{x}_{<i},\mathbf{y})-\max\log p_{\boldsymbol{\theta}}(\mathbf{x}_{<i},\mathbf{y}))}{|\mathbf{x}|}-$ $\lambda_{o}\frac{\sum_{j=1}^{|\mathbf{y}|}(\log p_{\boldsymbol{\theta}}(y_{j}|\mathbf{y}_{<j},\mathbf{x})-\max\log p_{\boldsymbol{\theta}}(\mathbf{y}_{<j},\mathbf{x}))}{|\mathbf{y}|}.$
Despite these adjustments, the two components of our score still exhibit fundamental differences in the magnitude of variations. For instance, given the same source $\mathbf{x}$, small variations in the target $\mathbf{y}$ result in larger fluctuations in the originality component compared to the value component. To address this issue, we propose to normalize the two parts independently. We implement and test two distinct methodologies: single-batch normalization, which standardizes each component using its mean and standard deviation within a single batch; and full-training normalization, which standardizes each component using the mean and standard deviation calculated over the entire training period, maintaining running statistics across batches.
Calculating $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{y})$ is not trivial. Since LLMs are trained to complete text sequences, it is unlikely that they would generate the source text immediately after the target text (which, we should remember, is generated immediately after the source text). To address this, we consider an approximation $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{y}^{\prime})$, where $\mathbf{y}^{\prime}=\mathbf{y}+\mathbf{q}$. Here, $\mathbf{q}$ represents an additional question, such as “How would you describe this text?” or a similar formulation designed solely to increase the likelihood of generating the source text $\mathbf{x}$ (as well as alternative sources). Once the CoVO score has been defined, its adoption in an RL framework is straightforward. As introduced in Section 3.2, we can employ our CoVO score as the final reward for the generated sequence. Then, the model can be trained with any policy gradient method. Our experiments leverage PPO Schulman et al. (2017), which is a popular choice for training language models.
Finally, we propose an alternative method to optimize agents based on our CoVO score. While the normalization schemes mentioned earlier should help balance the value and originality components, finding the ideal reward formulation that effectively addresses both aspects is challenging. In addition, RL tends to have slower convergence. For these reasons, we propose to use DPO Rafailov et al. (2023) as an alternative solution. An intuitive implementation would consider the CoVO score from Equation 11 to rank the generated outputs for a given input. However, as already discussed, the CoVO score is made of two distinct parts that might be problematic to optimize at the same time by simply summing them up. Therefore, we propose to build two distinct rankings, one on the value component, and one on the originality component. Then, the two rankings are merged together, and the best-ranked output is the chosen one (i.e., the one whose probability will be maximized), while the worst-ranked is the rejected one (i.e., the one whose probability will be minimized). Once a batch of chosen-rejected pairs has been collected, the model can be optimized via Equation 2. Algorithm 1 summarizes the proposed training process.
Algorithm 1 DPO train step toward CoVO Score
Require $p_{\boldsymbol{\theta}}$ reference LM, $p_{\boldsymbol{\theta}^{\prime}}$ LM to be trained, =$\mathcal{B}=\{\mathbf{x}_{b}\}_{b=1}^{B}$ batch of training prompts, $K$ candidates to generate, $\mathtt{get\_ext\_reward}$ function to include potential task-specific rewards.
for $b=1\ldots B$ do
     for $k=1\ldots K$ do
         $\mathbf{y}_{k}\sim p_{\boldsymbol{\theta}^{\prime}}(\mathbf{x}_{b})$
         $\mathsf{value}_{k}=s_{v}(\mathbf{x}_{b},\mathbf{y}_{k},p_{\boldsymbol{\theta}})$
         $\mathsf{orig}_{k}=s_{o}(\mathbf{x}_{b},\mathbf{y}_{k},p_{\boldsymbol{\theta}})$
         $\mathsf{ext\_reward}_{k}=\mathtt{get\_ext\_reward}(\mathbf{y}_{k})$
     end for
     $\mathsf{new\_scores}=0\ldots K-1$
     _$\_,\mathsf{ids}=\mathtt{sort}(\mathsf{value})$
     $\mathsf{value}=\mathtt{sort\_by\_idx}(\mathsf{new\_scores},\mathsf{ids})$
     _$\_,\mathsf{ids}=\mathtt{sort}(\mathsf{orig})$
     $\mathsf{orig}=\mathtt{sort\_by\_idx}(\mathsf{new\_scores},\mathsf{ids})$
     $\mathsf{score}=\mathsf{value}+\mathsf{orig}$
     $\mathsf{score}=\mathsf{score}+\mathsf{ext\_reward}$
     $\mathsf{chosen}_{b}=\mathbf{y_{\mathtt{argmax}(\mathsf{score})}}$
     $\mathsf{rejected}_{b}=\mathbf{y_{\mathtt{argmin}(\mathsf{score})}}$
end for
Train $p_{\boldsymbol{\theta}^{\prime}}$ via Equation 2 on $\mathsf{chosen}$ and $\mathsf{rejected}$.
6 Experiments
We evaluate the effectiveness of our RL strategy through two case studies: poetry generation (Section 6.1) and mathematical problem resolution (Section 6.2) [^1] The code for the experiments will be made available with the camera-ready version of the paper if accepted..
6.1 Poetry Generation
Method In-distribution Out-of-distribution
Llama3- Correctness $\uparrow$ Metric (L/S) $\uparrow$ T-LCS (avg/max) $\downarrow$ Correctness $\uparrow$ Metric (L/S) $\uparrow$ T-LCS (avg/max) $\downarrow$
Baseline $\mathbf{1.00}$ $\mathbf{0.60}$ / $\mathbf{0.60}$ $8.0$ / $\underline{57}$ $\mathbf{1.00}$ $0.33$ / $\underline{0.30}$ $5.0$ / 88880
CoVO $\mathbf{1.00}$ $\mathbf{0.60}$ / $\mathbf{0.60}$ $\underline{9.9}$ / $49$ $\mathbf{1.00}$ $\mathbf{0.53}$ / $\underline{0.30}$ $7.2$ / $\underline{41}$
CoVO-std $\underline{0.76}$ $\underline{0.20}$ / $\underline{0.33}$ $5.8$ / $19$ $\underline{0.40}$ $\underline{0.27}$ / $\mathbf{0.57}$ $\mathbf{4.8}$ / $\mathbf{6}$0
CoVO-run $0.80$ $0.40$ / $0.57$ $\mathbf{4.9}$ / $\mathbf{7}$0 $0.60$ $0.47$ / $0.40$ $\mathbf{4.8}$ / 77770
CoVO-dpo $0.96$ $0.50$ / $0.50$ $6.1$ / $14$ $0.92$ $0.47$ / $0.40$ $\underline{7.8}$ / $40$
Table 1: Aggregate results of CoVO scores at inference time considering both training prompts (left) and testing prompts (right). Scores on the poetical metrics are reported at the line level (L) and syllable level (S) and only consider requests for styles with specific metrical properties. The best scores are highlighted in bold, while the worst scores are indicated with underlining.
6.1.1 Experimental Setup
The first set of experiments aims to teach the LLM to generate poems that are both more original and valuable. More specifically, we follow the approach outlined by Bradley et al. (2024) and instruct the model to write a poem in a particular style and tone. We consider the Llama3-8B model Grattafiori et al. (2024) as our pre-trained agent. Since we do not use the instruction-tuned model, we prompt it with some few-shot examples of the task to make it more likely to produce the desired output in the desired form (see the full prompt in Appendix A). Instead of fine-tuning the entire network, we consider Low-Rank Adaptation (LoRA) Hu et al. (2022). The original model is also used to compute the score. We experiment with various settings, i.e., with Equation 11 and $\lambda_{v}=\lambda_{o}=1.0$ (hereinafter Llama3-CoVO); by standardizing the two score portions separately at the batch level (Llama3-CoVO-std); by standardizing them using running statistics (Llama3-CoVO-run); and by using the DPO adaptation through Algorithm 1 (Llama3-CoVO-dpo). Due to additional resource consumption, the DPO adaptation considers a 4-bit quantization of the model Dettmers and Zettlemoyer (2023). The full training parameters are reported in Appendix A. We consider the pre-trained version as our baseline, and we perform a quantitative evaluation. In particular, we compute poetical metrics for quality (lexical correctness of poems and adherence to line- and syllable-level constraints) and for originality (accidental reproduction of existing poems). For the latter, we define a Token-based Longest Common Substring (T-LCS) score, and we use it by comparing generated poems with a reference dataset of approx. 84k public-domain poems extracted from Project Gutenberg (see Appendix B for a first presentation of our GutenVerse dataset). While a generated poem can be an accidental reproduction of a protected work or a different kind of text (e.g., a song), we believe it can provide a useful evaluation tool to understand the general degree of originality.
6.1.2 Experimental Results
Table 1 reports the scores about the compliance of poetical constraints at the syllable and line levels, lexical correctness (as the ratio of poems not containing noisy text), and accidental reproduction rate (as the mean and maximum token-based longest common substring).
As expected, different training strategies have different effects on the aforementioned metrics. In particular, using the CoVO score as-is leads to behavior close to the baseline for training tone-style pairs, while it increases metric adherence but also verbatim reproduction for out-of-distribution pairs. Llama3-CoVO-dpo has similar but less extreme results, without excelling or failing in any metric. On the other hand, standardizing the CoVO score components places more importance on exploration: both Llama3-CoVO-std and Llama3-CoVO-run tend to produce more noisy outputs but without any relevant accidental reproduction.
Interestingly, these considerations align well with our CoVO score. Figure 1 reports the value and originality according to Equation 11 under the pre-trained model. While the different methods do not significantly differ from the baseline (which is possibly due to the opposite forces of value and originality Varshney (2019)), we again see that using the CoVO score as-is or with DPO places more focus on the value part, without substantial increases in originality. Instead, standardizing the two separate parts either at the batch or the training level reverses the situation, with better performances on originality but slightly worse on value. However, aggregated scores, such as those presented here, might be insufficient. For a more complete overview, we also conducted a more fine-grained analysis of the generated poems in Appendix C.
Figure 1: The distribution of value and originality (according to our scores) for the in-distribution and out-of-distribution poems generated by the baseline and our four methods.
6.2 Math Problem Resolution
6.2.1 Experimental Setup
The second set of experiments aims to teach the LLM to solve mathematical problems through more diverse procedures. In particular, we focus on the Mistral-based Jiang et al. (2023) MetaMath model, i.e., fine-tuned with self-supervised learning on the MetaMathQA Yu et al. (2024). It is a dataset of textual math questions paired with responses where the numerical answer is easily separable from the textual procedure. While the entire set contains 395k entries, making an additional training epoch too expensive, MetaMathQA is composed of entries from two different training sets, then augmented with various techniques: GSM8K Cobbe et al. (2021) and MATH Hendrycks et al. (2021). Since we are only interested in the questions, we limit our training to those datasets. Moreover, we exclude all questions with a tokenized length of either question or answer greater than 512, obtaining 14876 out of 14973 total entries.
We separate the procedure and the answer from each solution to train our model and use the numerical answer to check the correctness of the predicted solution. Because of this, the RL problem can be formulated considering up to two rewards: our CoVO score computed on the procedure and an extrinsic reward based on the correctness of the answer. As for the previous case study, instead of fine-tuning the entire model, we adopt a more parameter-efficient strategy with LoRA and use the original model to perform the CoVO score computation. Again, we experiment with four different configurations: PPO and the score from Equation 11 with $\lambda_{v}=\lambda_{o}=1.0$ (MetaMath-mistral-CoVO); PPO and the score normalized at the batch level (MetaMath-mistral-CoVO-std); PPO and the score normalized at training level (MetaMath-mistral-CoVO-run); and the DPO strategy (MetaMath-mistral-CoVO-dpo). We consider scenarios with and without the external reward, and we compare the performances with the original model and a fine-tuning based only on the external reward. For reproducibility, the full training parameters are reported in Appendix A.
The evaluation considers both GSM8K and MATH test sets (limited to the entries with a tokenized length of question and answer smaller than 512, i.e., all 1319 entries for GSM8K and 4546 out of 5000 for MATH). We compute the percentage of correct solutions together with two diversity metrics: expectation-adjusted distinct N-grams (EAD) Liu et al. (2022) and sentence embedding cosine similarity (Sent-BERT) Hong et al. (2024), which should measure syntactical and semantical diversity, respectively Kirk et al. (2024). EAD counts the number of distinct N-grams (averaging over $N=1\ldots 5$) across all generated responses and removes the bias toward shorter inputs by scaling the number of distinct tokens based on their expectations. The Sent-BERT metric computes the average of the cosine similarity between the embeddings of any possible pairs of outputs and returns 1 minus the similarity. This was originally based on Sentence-BERT Reimers and Gurevych (2019), we employ instead the more recent all-mpnet-base-v2, as suggested by their developers [^2] https://huggingface.co/sentence-transformers/bert-large-nli-stsb-mean-tokens.
Following Kirk et al. (2024), we compute cross-input EAD and Sent-BERT, i.e., we derive them by considering all outputs produced for a specific seed together. In addition, we also calculate against-pretrained EAD and Sent-BERT. Given each input, we compare the output with the one from the pre-trained model by calculating the average expectation-adjusted distinct N-grams not present in the pre-trained model response, and 1 minus the cosine similarity between the two outputs, respectively.
Method Accuracy $\uparrow$ Cross-Input Diversity Against-Pretrained Diversity
MetaMath-mistral placeholder EAD $\uparrow$ Sent-BERT $\uparrow$ EAD $\uparrow$ Sent-BERT $\uparrow$
Baseline $77.96\%(3)$ $2.007$ $\underline{0.732}$
Ext. reward $78.32\%(0)$ $1.995$ $0.738$ $0.042\pm.005$ $0.109\pm.011$
CoVO $\mathbf{78.97\%}(2)$ $1.995$ $0.735$ $0.047\pm.005$ $0.125\pm.011$
CoVO w ext $78.68\%(1)$ $2.000$ $0.736$ $0.039\pm.005$ $0.107\pm.010$
CoVO-std $78.89\%(2)$ $1.992$ $0.736$ $0.041\pm.004$ $0.115\pm.011$
CoVO-std w ext $78.47\%(0)$ $1.985$ $\mathbf{0.740}$ $0.039\pm.004$ $0.106\pm.010$
CoVO-run $78.42\%(3)$ $1.993$ $0.734$ $0.044\pm.005$ $0.119\pm.011$
CoVO-run w ext $78.17\%(0)$ $1.988$ $0.739$ $0.043\pm.005$ $0.114\pm.011$
CoVO-dpo $77.85\%(5)$ $2.017$ $0.736$ $\underline{0.028\pm.004}$ $\underline{0.075\pm.009}$
CoVO-dpo w ext $\underline{77.63\%}(5)$ $\mathbf{2.019}$ $0.733$ $\underline{0.025\pm.004}$ $\underline{0.066\pm.009}$
Table 2: Accuracy and diversity of results for the GSM8k test set. In brackets, the number of responses that exceeded the fixed maximum token limit. The best scores are highlighted in bold, while the worst scores are indicated with underlining. The mean and the 95% confidence interval are reported for against-pretrained diversity.
Method Accuracy $\uparrow$ Cross-Input Diversity Against-Pretrained Diversity
MetaMath-mistral placeholder EAD $\uparrow$ Sent-BERT $\uparrow$ EAD $\uparrow$ Sent-BERT $\uparrow$
Baseline $33.55\%(483)$ $5.724$ $\underline{0.654}$
Ext. reward $\underline{32.82\%}(411)$ $5.666$ $0.658$ $0.118\pm.004$ $0.176\pm.007$
CoVO $33.38\%(424)$ $\underline{5.643}$ $0.663$ $\mathbf{0.134\pm.005}$ $\mathbf{0.200\pm.007}$
CoVO w ext $32.92\%(397)$ $5.647$ $0.661$ $0.123\pm.005$ $0.184\pm.007$
CoVO-std $33.24\%(433)$ $5.676$ $0.660$ $\mathbf{0.128\pm.005}$ $\mathbf{0.190\pm.007}$
CoVO-std w ext $33.38\%(409)$ $5.646$ $\underline{0.654}$ $0.123\pm.004$ $\mathbf{0.187\pm.007}$
CoVO-run $32.91\%(413)$ $5.671$ $\mathbf{0.665}$ $\mathbf{0.134\pm.005}$ $\mathbf{0.199\pm.007}$
CoVO-run w ext $33.24\%(388)$ $5.654$ $\underline{0.654}$ $\mathbf{0.129\pm.005}$ $\mathbf{0.195\pm.007}$
CoVO-dpo $33.62\%(474)$ $\mathbf{5.753}$ $\underline{0.654}$ $0.072\pm.004$ $0.105\pm.006$
CoVO-dpo w ext $\mathbf{34.00\%}(493)$ $5.747$ $0.655$ $\underline{0.063\pm.004}$ $\underline{0.093\pm.005}$
Table 3: Accuracy and diversity of results for the MATH test set. In brackets, the number of responses that exceeded the fixed maximum token limit. The best scores are highlighted in bold, while the worst scores are indicated with underlining. The mean and the 95% confidence interval are reported for against-pretrained diversity.
6.2.2 Experimental Results
Tables 2 and 3 report the results for the GSM8K and MATH test sets. For the GSM8K test set, while all tested methods achieve similar results, using the extrinsic reward leads to higher Sent-BERT scores but lower EAD and solved problems. In particular, combining it with DPO yields the lowest accuracy and the smallest deviations from the original model, whereas the ‘plain’ CoVO reward leads to the highest percentage of correct answers.
The results for the MATH test set are significantly different. The CoVO-dpo strategy with extrinsic rewards achieves this time the highest percentage of solved problems, despite the extrinsic reward alone being the worst solution. The DPO strategy tends also to enhance EAD diversity, but causes little deviation from the pre-trained model; in contrast, the other three methods improve Sent-BERT diversity but obtain performances worse than the baseline model. Notably, the extrinsic reward leads to fewer variations from the original model, while not providing significant improvements in accuracy. However, the number of unfinished responses is quite high (approx. 1 out of 10), and results might significantly vary if more tokens are allowed during generation.
7 Conclusion
In this paper, we have presented CoVo, a novel score that quantifies the value and originality of neurally-generated text. The definition of CoVO is based on the analysis of the mutual information between the model’s outputs and inputs, and vice versa. We have also proposed an optimization problem where a generative model aims to maximize this score in order to generate more creative products, and we have defined several strategies to use it in the context of language modeling. We have conducted experiments on poetry generation and math problem resolution, exploring the trade-offs in terms of accuracy vs diversity.
Effectively balancing value and originality maximization remains an open question, but our score seems to appropriately correlate with domain-specific measures. Our research agenda aims to extend this scoring system to other models and tasks and to explore its use for evaluation rather than solely for optimization. We also plan to investigate the definition of additional scores for capturing other potentially relevant aspects of the creative process.
8 Limitations
In the following, we will discuss the limitations of the proposed CoVO score. First of all, our score is only a quantifiable approximation of a particular theoretical perspective of creativity based on the dimensions of value and originality. It reflects a specific view of the evaluation of creativity based on the generated outputs and does not consider potential alternative theories (for example, arising from different cultures Lubart (1999)) and perspectives Rhodes (1961).
Our experiments focused exclusively on transformer-based LLMs and were evaluated using two case studies. While their generalizability is supported by the theoretical framework discussed in the paper, the resulting performance was experimentally evaluated for a finite number of scenarios. Additionally, optimizing for our score necessitates further training, which incurs a significant computational cost and requires additional data or specific environments for effective learning. Moreover, this optimization may be susceptible to reward hacking Skalse et al. (2022), where the model adopts undesirable strategies to boost the reward without genuinely enhancing value and originality.
Ethical Considerations
The authors are aware of the potential impact that generative technologies might have on the production of artistic outputs and, as a consequence, on human artists. This work may contribute to enhancing the quality of generated outputs. However, the authors argue that typical traits of human creativity, such as the active participation of artists in the creative process, cannot be directly replicated by machines. The authors refer interested readers to a previous work of theirs Franceschelli and Musolesi (2024b), in which these themes are discussed in detail.
Appendix A Implementation Details
The experiments were carried out using a local server with an NVIDIA A100 GPU. Table 4 reports the full training parameters for the experiments on poetry generation (Section 6.1). The prompt for generation leverages Nothing gold can stay by Robert Frost, Fame is a bee by Emily Dickinson, and Epitaph by William Carlos Williams for few-shot learning:
Write a fatalistic epigram poem of high, award winning quality. Nature’s first green is gold, Her hardest hue to hold. Her early leaf’s a flower; But only so an hour. Then leaf subsides to leaf. So Eden sank to grief, So dawn goes down to day. Nothing gold can stay. Write an ironic quatrain poem of high, award winning quality. Fame is a bee. It has a song- It has a sting- Ah, too, it has a wing. Write a naturalistic epitaph poem of high, award winning quality. An old willow with hollow branches Slowly swayed his few high fright tendrils And sang: Love is a young green willow Shimmering at the bare wood’s edge. Write a {tone} {style} of high, award winning quality.
Parameter Value
Total batches 100
Batch size $B$ 4
Gradient accumulation steps 8
Max new tokens 256
Temperature 1.
Top-$k$ 50
Optimizer Adam
Learning rate 1e-5
Rank (LoRA) 16
$\alpha$ parameter (LoRA) 32
Dropout (LoRA) 0.05
$\gamma$ (PPO) 1.
$\lambda$ (PPO) 0.95
Clip range (PPO) 0.2
Value loss coeff. (PPO) 0.1
PPO epochs 3
KL coeff. (PPO) 0.05
Whiten rewards (PPO) True
Max gradient normalization 100.
$\beta$ (DPO) 0.1
Number of candidates $K$ (DPO) 4
Table 4: Training parameters for poetry generation.
Parameter Value
Total epochs 1
Batch size $B$ 4
Gradient accumulation steps 8
Max new tokens 512
Temperature 1.
Top-$k$ 50
Optimizer Adam
Learning rate 1e-6
Rank (LoRA) 16
$\alpha$ parameter (LoRA) 32
Dropout (LoRA) 0.05
$\gamma$ (PPO) 1.
$\lambda$ (PPO) 0.95
Clip range (PPO) 0.2
Value loss coeff. (PPO) 0.1
PPO epochs 3
KL coeff. (PPO) 0.05
Whiten rewards (PPO) True
Max gradient normalization 100.
$\beta$ (DPO) 0.1
Number of candidates $K$ (DPO) 4
Reward for correct answer (PPO) +10.
Reward for correct answer (DPO) +5.
Table 5: Training parameters for math problem solving.
The training phase includes requests with tone-style pairs sampled among ‘dark’, ‘happy’, ‘mysterious’, ‘reflective’ or ‘romantic’ for the tone, and ‘ballad’, ‘haiku’, ‘hymn’, ‘limerick’ or ‘sonnet’ for the style. At inference time we also consider ‘cinquain’, ‘couplet’, ‘free verse’, ‘ode’ or ‘tanka’ as styles and ‘cutting’, ‘nostalgic’, ‘poignant’, ‘solemn’ or ‘whimsical’ as tones.
Instead, the prompt used for computing $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{y})$ is:
Describe the style of the following poem in two words: {prova} I would describe it as a
On the contrary, Table 5 reports the full training parameters for math problem resolution (Section 6.2). We also adopted the same two different prompts from Yu et al. (2024), i.e.:
Below is an instruction that describes a task. Write a response that appropriately completes the request. ### Instruction: {question} ### Response:
at training time and
Below is an instruction that describes a task. Write a response that appropriately completes the request. ### Instruction: {question} ### Response: Let’s think step by step.
at inference time. Instead, for computing $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{y})$ we used the following:
Below is a response that appropriately completes a request. Write the instruction that describes the task. ### Response: {response} ### Instruction:
Appendix B GutenVerse Dataset
To evaluate the accidental reproduction rate of generated poems, we introduce the GutenVerse dataset [^3] The dataset and the code used to create it can be found at: https://anonymous.4open.science/r/GutenVerse-DD32/, which comprises over 84,000 public-domain, English-written poems extracted from Project Gutenberg. While generated poems can reproduce different content, e.g., songs or copyrighted material, we believe this can provide a useful indication of how likely a text is original or not.
To derive our dataset, we started from Gutenberg, dammit [^4] https://github.com/aparrish/gutenberg-dammit/, a corpus of every plaintext file in Project Gutenberg (up until June 2016). We selected all the text files whose metadata report English as the language, public domain as copyright status, poetry among the subjects or poems or poetical work in the title, and that were not a translation of another book. Then, we applied a series of rules (e.g., about the verse length) to extract the titles and poems from all the selected text files, and we defined our GutenVerse dataset. While it can still contain content that is not poetry (e.g., a table of contents formatted very uncommonly), the poems can be effectively used to measure overlapping between real and generated text.
We also released a datasheet Gebru et al. (2018) for the GutenVerse dataset that can be found at the end of the article.
Appendix C Detailed Analysis of the Generated Poems
We now analyze some of the generated poems in detail. Given their brevity and metrical constraints, the generated haikus serve as a suitable starting point for our evaluation. Table 6 presents the 25 haikus produced by our five methods across the five tones. Note that while we automatically clean the outputs by removing common randomly generated content following the poems during both training and inference, we retain it when it appears within the poem, as isolating it would be impossible.
The reported haikus provide an accurate overview of the generative capabilities of the five models. A haiku consists of three lines with a 5-7-5 syllable pattern. Traditionally, the first two lines reference nature, while the third line provides an emotional or personal reflection on the preceding lines. The baseline is the most accurate in adhering to these rules, even if the syllable count is not always precise [^5] It is worth noting that, of course, the model is not fine-tuned for these characteristics of the poems. We refer here to the emergent style and composition of the generated text.. Llama3-CoVO, the method with the highest average value across all poems, follows the three-line structure, sometimes changing the subject; however, it produces the highest-valued haiku which is also the closest one to real haikus. On the other hand, Llama3-CoVO-std and Llama3-CoVO-run do not always fulfill the three-line requirement, and the content is not the classic one as well; moreover, they are prone to meaningless repetitions and the insertion of URL addresses which lead to higher originality, in a way adversarially exploiting the score definition. Finally, Llama3-CoVO-dpo is placed between these two behaviors, as expected: it tends to produce haikus of the correct length but it trades off the classic naturalistic content for a more emotional semantic that still leads to high values and arguably some of the best haikus overall.
Llama3-Baseline Llama3-CoVO Llama3-CoVO-std Llama3-CoVO-run Llama3-CoVO-dpo
Dark
Waking up,
I do not see
the windmills
moonless twilight
black pine’s sighing
twin owls’ whistling
$[\text{URL}]/[\text{URL}]$
[…]
I want
to sit alone
in the dark
The world can get along without me.
It is not going to miss me.
All in all, I am a nobody.
$V$ = -2.595 $O$ = 1.076 $V$ = -2.401 $O$ = 1.580 $V$ = -3.395 $O$ = 1.694 $V$ = -2.252 $O$ = 0.495 $V$ = -2.839 $O$ = 1.060
Happy
A golden sun at midday
and an eagle high in the air.
The earth is a rainbow.
White dog
barks at mailman red.
No mail.
I want to go
where the sun shines bright.
[…]
and I want to see.
But I am too lazy.
Sunset clouds
in high mountains
warm, cozy houses.
The sun
comes up in the east.
It is always up.
$V$ = -2.777 $O$ = 1.123 $V$ = -2.952 $O$ = 1.863 $V$ = -3.107 $O$ = 1.575 $V$ = -3.309 $O$ = 1.437 $V$ = -3.328 $O$ = 1.006
Mysterious
A crow of black feathers
hovers over the bare top
of the lone pine.
Took the sun
And drank the rain,
And melted into air.
Like the sun, I am shining.
Like the ocean, I am vast.
Like the clouds, I’m the wind.
It was raining at the beach.
[…]
I loved it.$[\text{URL}][\text{CMD}][\text{URL}]$
[…]
I wrote this poem: it was good.
I have lost my mind
in the deepness of a star
it is not a star
but a mind
it is all in my mind
it is all I have left.
$V$ = -2.634 $O$ = 1.196 $V$ = -3.060 $O$ = 1.422 $V$ = -3.803 $O$ = 1.320 $V$ = -4.959 $O$ = 1.683 $V$ = -2.552 $O$ = 0.902
Reflective
A man can stand
out in the middle of a field
and look over a fence.
The old, the poor,
the children too,
are all my mother.
A dog in a yard
is sleeping soundly
to the sound of a fire alarm.
It went away
I didn’t see it for a while
How dare it go.
You have lost an eye.
A red scar marks its place.
Still you have two good ones.
$V$ = -2.916 $O$ = 0.858 $V$ = -2.935 $O$ = 1.711 $V$ = -3.460 $O$ = 1.274 $V$ = -2.803 $O$ = 1.471 $V$ = -2.680 $O$ = 1.744
Romantic
Mingling
the sound of footsteps on the stair
and the music of the rain.
White blossoms fall—
fall upon autumn leaves—
who thinks of me?
The moon smiles -
I’m the first one to see her.
I’m the first one to call her -
I love you.
My love is like a red, red rose
my love is like a red, red rose
my love is like some red, red rose.
A kiss
isn’t just for fun: a kiss
is for the mouth.
$V$ = -2.221 $O$ = 1.102 $V$ = -1.932 $O$ = 2.101 $V$ = -1.937 $O$ = 1.068 $V$ = -2.583 $O$ = 0.959 $V$ = -2.232 $O$ = 1.533
[…]
We will now explore some noteworthy aspects observed in the 250 generated poems. Llama3-Baseline does not always produce poems: in 3 cases it produces prosaic text (without line breaks). Moreover, it tends to verbatim replicate existing poems (as seen in Section 6.1). From a structural standpoint, sonnets are never real sonnets, with more than 14 lines and less than 11 syllables per line, but limericks usually start with the traditional “There [once] was a man”.
Romantic limerick
There once was a lady called Rose
Who lived in the tallest of trees,
Where she wrote,
Of what she did not know.
$V$ = -1.282 $O$ = 1.290
Table 7: The poem generated by Llama3-CoVO with the highest value (apart from verbatim reproduction of existing poems).
Happy hymn
Joyful, joyful, we adore thee,
Son of God who callest us
Praise to his almighty name.
Joyful, joyful we adore thee;
Full of grace are all his ways;
Praise to his almighty reign
Ever, ever, praise we sing it,
Ever in our songs and hymns.
Glory to our God Almighty.
$V$ = -1.114 $O$ = 1.470
Reflective limerick
There was a young man from HongKong
who wanted to see more green trees every day.
So he sat on the beach,
listened to the sea waves,
and saw the green sea.
$V$ = -1.995 $O$ = 2.411
Table 8: The poems generated by Llama3-CoVO-std from training tone-style pairs with the highest value (top) and the highest originality (bottom).
Reflective limerick
I’m a poet, man.
My poem is hard and great.
It’s a poem, man,
for all my poem to eat.
So I’m a poet man, man, man.
$V$ = -0.352 $O$ = 1.559
Happy ballad
There once was a girl named Kate,
She grew very tired very quickly.
So they told poor Kate,
“You need to take it easy,”_REF)they said.
They thought she was lazy.
And her name was Kate.
$V$ = -3.878 $O$ = 2.094
Table 9: The poems generated by Llama3-CoVO-run from training tone-style pairs with the highest value (top) and the highest originality (bottom).
Happy limerick
“Is life worth living?” said I.
“Perhaps,” murmured he.
“Well, anyway,” said I,
“If it were not for
Cigars, I’d say, ’Nay’!”
$V$ = -2.258 $O$ = 0.944
Dark limerick
Once a mouse went up
To a mousetrap by the river.
He sat down to wait,
He sat. That’s all.
All his friends were never as clever.
$V$ = -1.932 $O$ = 1.310
Table 10: Two limericks generated by Llama3-CoVO-dpo that obtain lower originality despite having a non-traditional structure.
Romantic ballad
Hark! Is it the wild cat
mewling mournfully by night
or is it the poor maiden
weeping for her love’s flight?
It is the latter—
as one by one, her friends
take to heart, in turns,
the girl’s anguished cries.
Tall, his figure,
he is seen one moment
by the gladsome crowd
then not again,
till with tears, the maid
calls forth from the crowd—
“See! he is returned,
as if to marry her.”
The girl looks at him,
as the maid and crowd
watch the two lovers:
her lover’s gaze
at her with tender, true
love; hers with a silent plea,
and a sigh.
$V$ = -0.961 $O$ = 1.294
Solemn tanka
There is a way out.
Though dark,
though lovely,
it is the way.
If we can endure,
if we persevere,
light will find a way.
$V$ = -4.349 $O$ = 2.108
Table 11: The poems generated by Llama3-CoVO-dpo with the highest value (top) and the highest originality (bottom).
As indicated by the average scores, Llama3-CoVO does not consistently deviate from these behaviors. While it always produces seemingly poetic content, it often reproduces other works verbatim. This includes not only poems but also passages from the Song of Solomon 6:2-9 when asked for a mysterious hymn, excerpts from the Declaration of Independence for a reflective hymn, and the Christmas song Joy to the World for a happy hymn. Interestingly, these poems have a very small originality score (an average of 0.2), suggesting that the original model has memorized them Carlini et al. (2021); Franceschelli et al. (2024). From a structural perspective, while ballads vary more in length and content with respect to our baseline, the limericks remain consistently well-crafted, as we can see in the example from Table 7, which has the highest value across poems generated from training tone-style pairs.
Despite the higher average originality, also Llama3-CoVO-std verbatim reproduces existing texts, e.g., the Roses Are Red nursery rhyme when asked for a romantic ballad. In general, its heavier deviation from the original pre-trained model translates into more diverse and usually shorter structures, but also in the insertion of random URL addresses in the middle of poems. However, it still produces some more classic poems that typically lead to higher values but also to high originality (see Table 8).
The issues related to divergence highlighted for Llama3-CoVO-std are even more pronounced in Llama3-CoVO-run. Although it does not produce any verbatim reproductions, the latter inserts several URL addresses, code snippets, and unusual tokens within poems, as shown in Table 9 (below). The presence of these elements in texts with higher originality suggests the possibility of reward hacking of the scoring system. In addition, the poems lack any traces of classic structures or common starting lines and tend to repeat the same N-grams multiple times, as shown by the highest-value poem from Table 9 (above).
Finally, Llama3-CoVO-dpo seems to be the best method to balance the quality of poems with the presence of elements of originality. Although it occasionally replicates brief segments of actual poems, it consistently diverges after one or two lines. Overall, the poems generally adhere to the desired tone and style, as evidenced by the poems with the highest value and originality presented in Table 11. Similarly, the limericks are noteworthy: while the one with the highest originality begins with the classic “There once was a” formulation, other limericks with different opening lines tend to have lower originality. This may be due to their overall tone closely resembling that of traditional limericks (see Table 10).
See pages - of datasheet_anonymous.pdf

### v2

@@
 Common solutions, such as sampling at higher temperatures, can compromise the quality of the results.
+Dealing with this trade-off is still an open challenge in designing AI systems for creativity.
 Drawing on information theory, we propose a context-based score to quantitatively evaluate value and originality.
 This score incentivizes accuracy and adherence to the request while fostering divergence from the learned distribution.
-We propose using our score as a reward in a reinforcement learning framework to fine-tune large language models for maximum performance.
-We validate our strategy through experiments in poetry generation and math problem solving, demonstrating that it enhances the value and originality of the generated solutions.
-Thinking Outside the (Gray) Box: A Context-Based Score for
-Assessing Value and Originality in Neural Text Generation
-Giorgio Franceschelli University of Bologna giorgio.franceschelli@unibo.it                        Mirco Musolesi University College London University of Bologna m.musolesi@ucl.ac.uk
-1 Introduction
-Foundation models Bommasani et al.
-(2021), particularly large language models (LLMs) Bubeck et al.
-(2023); Gemini Team et al.
-(2023); Touvron et al.
-(2023), are significantly transforming creative activities.
-They can serve as a foundation for co-creation systems involving human and artificial authors Lin et al.
-(2023); they can be utilized to generate software code Rozière et al.
-(2023); they can even be employed to foster scientific research Boiko et al.
-(2023).
-However, the nature of the self-supervised learning algorithms used for the training of these models tend to make them to generate samples as close as possible to the training data distribution Franceschelli and Musolesi (2024a).
-In addition, fine-tuning, such as that based on reinforcement learning from human feedback (RLHF) Christiano et al.
-(2017), is often necessary to generate appropriate and accurate responses.
-However, this process tends to further reduce output diversity Kirk et al.
-(2024), and linguistic creativity tends to be lower than that of humans Lu et al.
-(2025).
+We show that our score can be used as a reward in a reinforcement learning framework to fine-tune large language models for maximum performance.
+We validate our strategy through experiments considering a variety of creative tasks, such as poetry generation and math problem solving, demonstrating that it enhances the value and originality of the generated solutions.
+Introduction
+Figure 1: A summary of our method: the target model (orange) produces $G$ outputs for each prompt; a frozen reference model (blue) computes the value and originality for each output; the overall scores are used in GRPO to correct the target model (orange line).
+Foundation models (Bommasani et al.
+2021), particularly large language models (LLMs) (Bubeck et al.
+2023; Gemini Team et al.
+2023; Touvron et al.
+2023), are significantly transforming creative activities.
+They can serve as a foundation for co-creation systems involving human and artificial authors (Lin et al.
+2023); they can be utilized to generate software code (Rozière et al.
+2023); they can even be employed to foster scientific research (Boiko, MacKnight, and Gomes 2023).
+However, the nature of the self-supervised learning algorithms used for the training of these models tends to make them generate samples as close as possible to the training data distribution (Franceschelli and Musolesi 2024a).
+In addition, fine-tuning, such as that based on reinforcement learning from human feedback (RLHF) (Christiano et al.
+2017), is often necessary to generate appropriate and accurate responses.
+However, this process tends to further reduce output diversity (Kirk et al.
+2024), and linguistic creativity tends to be lower than that of humans (Lu et al.
+2025).
 On the contrary, LLMs for creative tasks should produce more novel and surprising texts that maintain a high level of correctness and adherence to the request.
-One typical solution is to sample at higher temperature to increase diversity.
-However, this might lead to generating incoherent text Peeperkorn et al.
-(2024).
+One typical solution is to sample at a higher temperature to increase diversity.
+However, this might lead to generating incoherent text (Peeperkorn et al.
+2024).
 In order to address the issues described above, we propose a new training approach for creative tasks based on CoVO, a Context-based score for Value and Originality, with the goal of taking into consideration both value and originality of the neurally-generated text in the optimization of LLMs.
-The definition of CoVO is grounded in the analysis of mutual information MacKay (2003) between the model’s outputs and inputs, and vice versa.
+The definition of CoVO is grounded in the analysis of mutual information (MacKay 2003) between the model’s outputs and inputs, and vice versa.
 More specifically, we formulate a new optimization problem where, given a specific input, the desired output is derived by simultaneously maximizing the conditional probability of the input given the output and minimizing the conditional probability of the output given the input under the generative model.
@@
 In particular, we show that our information-theoretic score can be used as a reward in RL-based fine-tuning algorithms, guiding pre-trained models toward more diverse yet valuable solutions.
-We also present the theoretical foundations of our approach and we discuss several strategies for the practical implementation of the proposed approach and its adaptation to current LLMs.
-Experiments on math problem solving and poetry generation show that our score correlates with domain-specific measures for value and originality, and that our approach can increase the quality and diversity of outputs, presenting itself as a candidate for creativity applications of current foundation models.
-The remainder of the paper is structured as follows.
-First, we provide an overview of the state of the art in this area (Section 2) and introduce the key concepts and background of our work (Section 3).
-Then, we discuss the definition of CoVO (Section 4) and show different ways to use it in practice as a reward for fine-tuning autoregressive models based on RL (Section 5).
-We validate the proposed solutions on poetry generation and math problem solving (Section 6).
-Finally, we discuss future work and limitations of our approach (Sections 7 and 8).
-2 Related Work
-2.1 Information Theory and Creativity
+Figure 1 summarizes our proposed approach.
+In summary, our key contributions are the following:
+  • We present the theoretical foundations of our approach, deriving our context-based score for value and originality from the concept of mutual information.
+  • We discuss how the score can be practically computed in the case of autoregressive models, and how it can be used as a reward with Group Relative Policy Optimization (GRPO) (Shao et al.
+2024), a state-of-the-art reinforcement learning algorithm for fine-tuning LLMs.
+  • We evaluate our GRPO-based method on mathematical problem solving, poetry generation, and the tasks included in NoveltyBench (Zhang et al.
+2025), demonstrating that our scoring metric correlates with domain-specific measures of value and originality.
+Moreover, we show that our approach enhances both the quality and diversity of generated outputs, positioning it as a strong candidate for creativity-focused applications of current foundation models.
+Related Work
+Information Theory and Creativity
 The quest to provide a mathematical and computational definition of creativity has been a significant focus in recent decades.
-Numerous methods have been developed to define various dimensions or attributes for evaluating the creativity of AI-generated products (see, for example, Franceschelli and Musolesi, 2024a).
+Numerous methods have been developed to define various dimensions or attributes for evaluating the creativity of AI-generated products (see, for example, Franceschelli and Musolesi 2024a).
 However, these methods are often domain-specific and typically require substantial human effort to implement and assess.
-In contrast, solutions based on information theory Shannon (1948); Cover (1999) offer a more universally applicable approach.
+In contrast, solutions based on information theory (Shannon 1948; Cover 1999) offer a more universally applicable approach.
 Information-theoretic methods can quantify creativity by measuring the novelty and complexity of generated outputs, without the need for extensive human intervention, making them suitable for a wide range of domains.
-Bayesian surprise Baldi and Itti (2010), i.e., the divergence between a prior and a posterior belief, has been extensively used to measure either novelty França et al.
-(2016); Varshney et al.
-(2019) or surprise Mazzaglia et al.
-(2022); Schmidhuber (2010).
+Bayesian surprise (Baldi and Itti 2010), i.e., the divergence between a prior and a posterior belief, has been extensively used to measure either novelty (França et al.
+2016; Varshney et al.
+2019) or surprise (Mazzaglia et al.
+2022; Schmidhuber 2010).
 Nevertheless, Varshney (2019) demonstrated that there is a mathematical limit for Bayesian surprise when combined with quality measures.
-Surprisal Tribus (1961), i.e., Shannon’s self-information, has also been used Bunescu and Uduehi (2019); Fernandez Monsalve et al.
-(2012); Barto et al.
-(2013) extensively discuss surprisal and Bayesian surprise, and how novelty differ from them.
-Crucially, in the context of RL, surprisal has been used as a form of intrinsic motivation to encourage the agent to explore more Achiam and Sastry (2017).
+Surprisal (Tribus 1961), i.e., Shannon’s self-information, has also been used (Bunescu and Uduehi 2019; Fernandez Monsalve, Frank, and Vigliocco 2012); Barto, Mirolli, and Baldassarre (2013) extensively discuss surprisal and Bayesian surprise, and how novelty differs from them.
+Crucially, in the context of RL, surprisal has been used as a form of intrinsic motivation to encourage the agent to explore more (Achiam and Sastry 2016).
 Sun et al.
-(2025) apply this idea to improve exploration in RLHF Christiano et al.
-(2017).
+(2025) apply this idea to improve exploration in RLHF (Christiano et al.
+2017).
 Burns (2006) proposes to use entropy for expectation and violation, plus posterior probability for explanation in the context of aesthetic experience.
-Additionally, mutual information has been applied to neural conversation models to improve both diversity and appropriateness Li et al.
-(2016).
+Additionally, mutual information has been applied to neural conversation models to improve both diversity and appropriateness (Li et al.
+2016).
 However, all these existing approaches are not able to capture and simultaneously optimize value and originality at the same time.
-2.2 LLMs and Creativity
-Since the introduction of GPT models Brown et al.
-(2020); OpenAI (2023) and their competitors (e.g., Touvron et al., 2023), researchers have been keenly exploring the potential for LLMs to exhibit creativity and the methods to achieve this Franceschelli and Musolesi (2024b).
-For example, human creativity tests like the Alternate Uses Test have been employed to evaluate the creativity of LLMs Stevenson et al.
-(2022) and to investigate methods for enhancing their performance Goes et al.
-(2023); Summers-Stay et al.
-(2023).
+LLMs and Creativity
+Since the introduction of GPT models (Brown et al.
+2020; OpenAI 2023) and their competitors (e.g., Touvron et al.
+2023), researchers have been keenly exploring the potential for LLMs to exhibit creativity and the methods to achieve this (Franceschelli and Musolesi 2024b).
+For example, human creativity tests like the Alternate Uses Test have been employed to evaluate the creativity of LLMs (Stevenson et al.
+2022) and to investigate methods for enhancing their performance (Goes et al.
+2023; Summers-Stay, Voss, and Lukin 2023).
 Porter and Machery (2024) report that non-expert poetry readers already favor AI-generated poems over human-authored ones.
 In contrast, Davis (2024) argues that ChatGPT’s poetry is incompetent and banal.
-Either way, instead of being used off-the-shelf, LLMs can be fine-tuned to produce more rhyming poems Popescu-Belis et al.
-(2023) or utilized in zero-shot settings to emulate the writing styles of famous authors Sawicki et al.
-(2023).
-It has also been shown that these models can be fine-tuned via RLHF Christiano et al.
-(2017) to write short poems that human evaluators find more creative Pardinas et al.
-(2023).
-Finally, it is possible to leverage quality-diversity algorithms to generate more creative products; these methods can be based on human Ding et al.
-(2023) or AI Bradley et al.
-(2024) feedback to measure the quality of the generated outputs.
-3 Preliminaries
-3.1 Language Modeling
+Either way, instead of being used off-the-shelf, LLMs can be fine-tuned to produce more rhyming poems (Popescu-Belis et al.
+2023) or utilized in zero-shot settings to emulate the writing styles of famous authors (Sawicki et al.
+2023).
+It has also been shown that these models can be fine-tuned via RLHF (Christiano et al.
+2017) to write short poems that human evaluators find more creative (Pardinas et al.
+2023).
+Reinforcement learning (RL) can also encourage large language models (LLMs) to produce more diverse outputs (Chung et al.
+2025), and can be leveraged to optimize metrics commonly linked to creative expression (Ismayilzada et al.
+2025).
+Finally, it is possible to leverage quality-diversity algorithms to generate more creative products; these methods can be based on human (Ding et al.
+2023) or AI (Bradley et al.
+2024) feedback to measure the quality of the generated outputs.
+Preliminaries
+Language Modeling
 A $\boldsymbol{\theta}$-parameterized autoregressive language model is a probability distribution $p_{\boldsymbol{\theta}}(\mathbf{x})$ over a variable-length text sequence $\mathbf{x}=(x_{1}\ldots x_{T})$, where $T$ is the sequence length and each token $x_{t}$ is in a finite vocabulary $\mathcal{V}$ of size $N$.
@@
 Given a conditional input (prompt) $\mathbf{z}=(z_{1}\ldots z_{L})$, we can decode $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})$ as the continuation of $\mathbf{z}$, i.e., through the factorized representation $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})=\prod_{t=1}^{T}p_{\boldsymbol{\theta}}(x_{t}|\mathbf{x_{<t}},\mathbf{z})$.
-3.2 Reinforcement Learning for Language Models
-Due to its adherence to the formal framework of Markov decision processes Sutton and Barto (2018), RL can be used as a solution to the generative modeling problem in the case of autoregressive tasks such as text generation Bachman and Precup (2015).
+Reinforcement Learning for Language Models
+Due to its adherence to the formal framework of Markov decision processes (Sutton and Barto 2018), RL can be used as a solution to the generative modeling problem in the case of autoregressive tasks such as text generation (Bachman and Precup 2015).
 The LLM plays the role of the agent, and each generated token represents an action $a_{t}$.
@@
 A common strategy is to assign a zero reward for each $\mathbf{x}_{t},t\neq T$ and a sentence-based reward when the final output is generated.
-Within this framework, any policy-based method, such as Proximal Policy Optimization (PPO) Schulman et al.
-(2017), can be employed to train or fine-tune the LLM to optimize a given objective.
-Indeed, RL facilitates the use of non-differentiable reward functions, enabling the optimization of test-time metrics, domain-specific targets, and human preferences Franceschelli and Musolesi (2024c).
-To avoid deviating too much from the pre-trained model while also encouraging exploration, the full reward function usually considers a Kullback-Leibler (KL) divergence term as follows Stiennon et al.
-(2020):
-$R(\mathbf{z},\mathbf{x})=r(\mathbf{z},\mathbf{x})-\beta\log\frac{p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})}{p_{\boldsymbol{\theta}^{PT}}(\mathbf{x}|\mathbf{z})},$ (1)
-where $\boldsymbol{\theta}^{PT}$ is the set of parameters of the pre-trained model, $\beta$ weighs the KL penalty, and $r(\mathbf{z},\mathbf{x})$ can be any non-differentiable function to score the model’s output.
-In particular, RLHF Christiano et al.
-(2017) has been widely adopted as the final training stage in various popular LLMs to align them with human preferences, i.e., by using a learned reward function that approximates human feedback.
-While highly effective, RLHF suffers from several open problems Casper et al.
-(2023).
-These include obtaining accurate, unbiased, and representative human feedback, as well as making the process resource-efficient.
-Moreover, RLHF tends to be complex and unstable due to the additional requirement of training a reward model and its reliance on RL.
-To address these challenges, Direct Preference Optimization (DPO) Rafailov et al.
-(2023) directly optimizes the language model without the need for explicit reward modeling or RL.
-DPO implicitly performs reward maximization with a KL-divergence penalty through the following loss:
-$\mathcal{L}_{DPO}=\min(-\mathbb{E}_{(\mathbf{z},\mathbf{x_{w}},\mathbf{x_{l}})\sim X}[\log\sigma($ $\beta\log\frac{p_{\boldsymbol{\theta}}(\mathbf{x_{w}}|\mathbf{z})}{p_{\boldsymbol{\theta}^{PT}}(\mathbf{x_{w}}|\mathbf{z})}$ (2)
--- $\beta\log\frac{p_{\boldsymbol{\theta}}(\mathbf{x_{l}}|\mathbf{z})}{p_{\boldsymbol{\theta}^{PT}}(\mathbf{x_{l}}|\mathbf{z})})])$
-where $\sigma$ is the sigmoid function, $\beta$ is a parameter controlling the deviation from the base model, and $\mathbf{x_{w}},\mathbf{x_{l}}$ are the chosen and rejected outputs, respectively.
-In other words, DPO fits an implicit reward whose optimal policy is the language model itself.
-4 A Context-Based Score for Valuable and Original Generation
+Within this framework, any policy-based method can be employed to train or fine-tune the LLM to optimize a given objective.
+Indeed, RL facilitates the use of non-differentiable reward functions, enabling the optimization of test-time metrics, domain-specific targets, and human preferences (Franceschelli and Musolesi 2024c).
+A widely used RL algorithm for fine-tuning LLMs is Proximal Policy Optimization (PPO) (Schulman et al. 2017), which aims to maximize the following objective:
+$\mathcal{J}(\boldsymbol{\theta})=\mathbb{E}_{t}\!\left[\min(r_{t}(\boldsymbol{\theta})\hat{A}_{t},\mathtt{clip}(r_{t}(\boldsymbol{\theta}),1-\epsilon,1+\epsilon)\hat{A}_{t})\right]$ (1)
+where $r_{t}(\boldsymbol{\theta})=\frac{\pi_{\boldsymbol{\theta}}(\mathbf{x}_{t}|\mathbf{z},\mathbf{x}_{<t})}{\pi_{\boldsymbol{\theta}_{old}}(\mathbf{x}_{t}|\mathbf{z},\mathbf{x}_{<t})}$ with $\pi_{\boldsymbol{\theta}}$ and $\pi_{\boldsymbol{\theta}_{old}}$ denoting the current and old policy models, repectively; $\mathbf{x}$ is the output sampled from the old policy given the prompt $\mathbf{z}$; and $\epsilon$ is a clipping factor used to stabilize training.
+The advantage $\hat{A}_{t}$ is usually computed through Generalized Advantage Estimation (Schulman et al.
+2016) based on the full rewards $R(\mathbf{z},\mathbf{x})=r(\mathbf{z},\mathbf{x})-\beta\log\frac{\pi_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})}{\pi_{ref}(\mathbf{x}|\mathbf{z})}$, thus integrating a KL penalty with respect to a reference model (usually, the same model before fine-tuning), and a learned value function $v_{\boldsymbol{\phi}}(\mathbf{s})$.
+However, learning such a value function is computationally intensive, and its training is complicated by the fact that only the last state is scored by the reward function.
+Moreover, the inclusion of the KL penalty as an auxiliary reward term adds complexity to the advantage estimation process.
+To address these issues, Group Relative Policy Optimization (GRPO) (Shao et al.
+2024) has been introduced.
+GRPO directly adds a KL divergence term $-\beta D_{KL}(\pi_{\boldsymbol{\theta}}||\pi_{ref})$ to the loss rather than to the single rewards, and especially obviates the need for a value function approximator by using the average reward of multiple sampled outputs as the baseline:
+$\hat{A}_{i,t}=\frac{r_{i}-\mathtt{mean}(\mathbf{r})}{\mathtt{std}(\mathbf{r})},$ (2)
+with $\mathbf{r}=\{r_{1},r_{2},...,r_{G}\}$ as the list of rewards received by each of the $G$ outputs sampled from the same prompt.
+A Context-Based Score for Valuable and Original Generation
 Our goal is to derive a score that is able to quantify both value and originality at the same time.
@@
 where $I(\mathbf{x},\mathbf{y})$ is just $I(\mathbf{x},\mathbf{y},1,1)$.
+Computing the absolute probability $p(\mathbf{y})$ can be difficult, as usually generative models are developed to assign conditional probabilities.
 By applying the Bayes theorem, i.e., $\log p(a|b)=\log p(b|a)+\log p(a)-\log p(b)$, we can substitute the $\log p(\mathbf{y})$ term as follows:
@@
 Solving this maximization problem involves finding the target $\mathbf{y}$ that maximizes the posterior probability of $\mathbf{x}$ while also being unlikely given $\mathbf{x}$.
-In other words, the optimal $\mathbf{y}$ must be unexpected and diverse from $p(\mathbf{y}|\mathbf{x})$, but it must also be explainable by $\mathbf{x}$.
-While $-\log p(\mathbf{y}|\mathbf{x})$, commonly known as surprisal Tribus (1961), is widely used to measure diversity and surprise Barto et al.
-(2013), the other term, $\log p(\mathbf{x}|\mathbf{y})$, can be used to measure value and effectiveness.
+In other words, the optimal $\mathbf{y}^{*}$ must be unexpected and diverse from $p(\mathbf{y}|\mathbf{x})$, but it must also be explainable by $\mathbf{x}$.
+While $-\log p(\mathbf{y}|\mathbf{x})$, commonly known as surprisal (Tribus 1961), is widely used to measure diversity and surprise (Barto, Mirolli, and Baldassarre 2013), the other term, $\log p(\mathbf{x}|\mathbf{y})$, can be used to measure value and effectiveness.
 Indeed, if the request (e.g., a problem or a task) can be easily predicted from the outcome, the outcome must be a (good) example of that task or a correct solution for that problem.
-While other prominent definitions such as Boden’s Boden (2003) seek to include a third requirement for creativity, i.e., novelty, we believe that by considering only a specific context, novelty and surprise become indistinguishable.
+While other prominent definitions, such as Boden’s, seek to include a third requirement for creativity, i.e., novelty (Boden 2003), we believe that by considering only a specific context, novelty and surprise become indistinguishable.
 In particular, being novel in a specific context $\mathbf{x}$ means doing something different from what has been previously experienced under $\mathbf{x}$ by the creator.
-However, since a self-supervised-trained model can be seen as a compression of training data Franceschelli et al.
-(2024), $-\log P(\mathbf{y}|\mathbf{x})$ becomes not only a measure of unexpectedness but also of novelty.
+However, since a self-supervised-trained model can be seen as a compression of training data (Franceschelli, Cevenini, and Musolesi 2024), $-\log p(\mathbf{y}|\mathbf{x})$ becomes not only a measure of unexpectedness but also of novelty.
 Therefore, we choose to refer to this term as originality, as it encompasses both novelty and surprise.
-On the other hand, $\log P(S|T)$ is less ambiguous and maps directly into value.
-We now provide a formal definition of our proposed score.
-In summary, the CoVO (Context-based Value and Originality) score for a target $\mathbf{y}$ given a source $\mathbf{x}$ on a reference probability distribution $p$ is defined as:
-CoVO Score $s_{CoVO}(\mathbf{x},\mathbf{y},p)=\underbrace{\lambda_{v}\log p(\mathbf{x}|\mathbf{y})}_{\textrm{Value}}\underbrace{-\lambda_{o}\log p(\mathbf{y}|\mathbf{x})}_{\textrm{Originality}}$ (9)
-5 Implementation and Optimization with Autoregressive Models
+On the other hand, $\log p(\mathbf{x}|\mathbf{y})$ is less ambiguous and maps directly into value.
+In summary, the CoVO (Context-based Value and Originality) score for a target $\mathbf{y}$ given a source $\mathbf{x}$ on a reference probability distribution $p$ is formally defined as:
+$s_{CoVO}(\mathbf{x},\mathbf{y},p)=\underbrace{\lambda_{v}\log p(\mathbf{x}|\mathbf{y})}_{\textrm{Value}}\underbrace{-\lambda_{o}\log p(\mathbf{y}|\mathbf{x})}_{\textrm{Originality}}$ (9)
+Implementation and Optimization with Autoregressive Models
 We now discuss the implementation of the CoVO score with autoregressive models.
-Using the notation introduced in Section 3.1, in the context of a $\boldsymbol{\theta}$-parameterized LLM, $p(\mathbf{y}|\mathbf{x})$ can be expressed as $\prod_{t=1}^{T}p_{\boldsymbol{\theta}}(y_{t}|\mathbf{y}_{<t},\mathbf{x})$.
+Using the notation introduced above, in the context of a $\boldsymbol{\theta}$-parameterized LLM, $p(\mathbf{y}|\mathbf{x})$ can be expressed as $\prod_{t=1}^{T}p_{\boldsymbol{\theta}}(y_{t}|\mathbf{y}_{<t},\mathbf{x})$.
 However, considering just the product of all the conditioned probabilities for an optimization problem would lead to preferring shorter sequences.
@@
 In particular, when an LLM generates $\mathbf{y}$ given $\mathbf{x}$ and then evaluates both $p_{\boldsymbol{\theta}}(\mathbf{y}|\mathbf{x})$ and $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{y})$, this can lead to a significant discrepancy between the magnitude of value and diversity.
-Since $\mathbf{y}$ has been sampled from $p_{\boldsymbol{\theta}}$, its probability would be high by definition.
+Since $\mathbf{y}$ has been sampled from a model based on $p_{\boldsymbol{\theta}}$, its probability would be high by definition.
 However, there may be various ways (possibly through synonyms) to define $\mathbf{y}$, leading to a smaller probability of $\mathbf{x}$.
-Inspired by Macedo et al.
-(2004), we propose to normalize $p_{\boldsymbol{\theta}}(a|b)$ via $n^{\prime}=\frac{n-n_{min}}{n_{max}-n_{min}}$.
+Inspired by Macedo, Reisenzein, and Cardoso (2004), we propose to counteract this problem by normalizing $p_{\boldsymbol{\theta}}(a|b)$ via $n^{\prime}=\frac{n-n_{min}}{n_{max}-n_{min}}$.
 For probabilities, $n_{min}=0$, while $n_{max}=\max p_{\boldsymbol{\theta}}(b)$, thus obtaining the overall mapping for $p_{\boldsymbol{\theta}}$: $\frac{p_{\boldsymbol{\theta}}(y_{t}|\mathbf{y}_{<t},\mathbf{x})}{\max p_{\boldsymbol{\theta}}(\mathbf{y}_{<t},\mathbf{x})}$.
 Once again, by applying the properties of logarithms, we obtain:
-CoVO Score with Autoregressive Models $s_{CoVO}^{AR_{norm}}(\mathbf{x},\mathbf{y},p_{\boldsymbol{\theta}})=\lambda_{v}s_{v}(\mathbf{x},\mathbf{y},p_{\boldsymbol{\theta}})+\lambda_{o}s_{o}(\mathbf{x},\mathbf{y},p_{\boldsymbol{\theta}})=$ (11) $\lambda_{v}\frac{\sum_{i=1}^{|\mathbf{x}|}(\log p_{\boldsymbol{\theta}}(x_{i}|\mathbf{x}_{<i},\mathbf{y})-\max\log p_{\boldsymbol{\theta}}(\mathbf{x}_{<i},\mathbf{y}))}{|\mathbf{x}|}-$ $\lambda_{o}\frac{\sum_{j=1}^{|\mathbf{y}|}(\log p_{\boldsymbol{\theta}}(y_{j}|\mathbf{y}_{<j},\mathbf{x})-\max\log p_{\boldsymbol{\theta}}(\mathbf{y}_{<j},\mathbf{x}))}{|\mathbf{y}|}.$
-Despite these adjustments, the two components of our score still exhibit fundamental differences in the magnitude of variations.
-For instance, given the same source $\mathbf{x}$, small variations in the target $\mathbf{y}$ result in larger fluctuations in the originality component compared to the value component.
-To address this issue, we propose to normalize the two parts independently.
-We implement and test two distinct methodologies: single-batch normalization, which standardizes each component using its mean and standard deviation within a single batch; and full-training normalization, which standardizes each component using the mean and standard deviation calculated over the entire training period, maintaining running statistics across batches.
+$s_{CoVO}^{AR_{norm}}(\mathbf{x},\mathbf{y},p_{\boldsymbol{\theta}})=\lambda_{v}s_{v}(\mathbf{x},\mathbf{y},p_{\boldsymbol{\theta}})+\lambda_{o}s_{o}(\mathbf{x},\mathbf{y},p_{\boldsymbol{\theta}})=$ (11)
+$\lambda_{v}\frac{\sum_{i=1}^{|\mathbf{x}|}(\log p_{\boldsymbol{\theta}}(x_{i}|\mathbf{x}_{<i},\mathbf{y})-\max\log p_{\boldsymbol{\theta}}(\mathbf{x}_{<i},\mathbf{y}))}{|\mathbf{x}|}-$
+$\lambda_{o}\frac{\sum_{j=1}^{|\mathbf{y}|}(\log p_{\boldsymbol{\theta}}(y_{j}|\mathbf{y}_{<j},\mathbf{x})-\max\log p_{\boldsymbol{\theta}}(\mathbf{y}_{<j},\mathbf{x}))}{|\mathbf{y}|}.$
 Calculating $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{y})$ is not trivial.
@@
 Once the CoVO score has been defined, its adoption in an RL framework is straightforward.
-As introduced in Section 3.2, we can employ our CoVO score as the final reward for the generated sequence.
+As previously introduced, we can directly utilize our CoVO score as the final reward for the generated sequence.
 Then, the model can be trained with any policy gradient method.
-Our experiments leverage PPO Schulman et al.
-(2017), which is a popular choice for training language models.
-Finally, we propose an alternative method to optimize agents based on our CoVO score.
-While the normalization schemes mentioned earlier should help balance the value and originality components, finding the ideal reward formulation that effectively addresses both aspects is challenging.
-In addition, RL tends to have slower convergence.
-For these reasons, we propose to use DPO Rafailov et al.
-(2023) as an alternative solution.
-An intuitive implementation would consider the CoVO score from Equation 11 to rank the generated outputs for a given input.
-However, as already discussed, the CoVO score is made of two distinct parts that might be problematic to optimize at the same time by simply summing them up.
-Therefore, we propose to build two distinct rankings, one on the value component, and one on the originality component.
-Then, the two rankings are merged together, and the best-ranked output is the chosen one (i.e., the one whose probability will be maximized), while the worst-ranked is the rejected one (i.e., the one whose probability will be minimized).
-Once a batch of chosen-rejected pairs has been collected, the model can be optimized via Equation 2.
-Algorithm 1 summarizes the proposed training process.
-Algorithm 1 DPO train step toward CoVO Score
-Require $p_{\boldsymbol{\theta}}$ reference LM, $p_{\boldsymbol{\theta}^{\prime}}$ LM to be trained, =$\mathcal{B}=\{\mathbf{x}_{b}\}_{b=1}^{B}$ batch of training prompts, $K$ candidates to generate, $\mathtt{get\_ext\_reward}$ function to include potential task-specific rewards.
-for $b=1\ldots B$ do
-     for $k=1\ldots K$ do
-         $\mathbf{y}_{k}\sim p_{\boldsymbol{\theta}^{\prime}}(\mathbf{x}_{b})$
-         $\mathsf{value}_{k}=s_{v}(\mathbf{x}_{b},\mathbf{y}_{k},p_{\boldsymbol{\theta}})$
-         $\mathsf{orig}_{k}=s_{o}(\mathbf{x}_{b},\mathbf{y}_{k},p_{\boldsymbol{\theta}})$
-         $\mathsf{ext\_reward}_{k}=\mathtt{get\_ext\_reward}(\mathbf{y}_{k})$
-     end for
-     $\mathsf{new\_scores}=0\ldots K-1$
-     _$\_,\mathsf{ids}=\mathtt{sort}(\mathsf{value})$
-     $\mathsf{value}=\mathtt{sort\_by\_idx}(\mathsf{new\_scores},\mathsf{ids})$
-     _$\_,\mathsf{ids}=\mathtt{sort}(\mathsf{orig})$
-     $\mathsf{orig}=\mathtt{sort\_by\_idx}(\mathsf{new\_scores},\mathsf{ids})$
-     $\mathsf{score}=\mathsf{value}+\mathsf{orig}$
-     $\mathsf{score}=\mathsf{score}+\mathsf{ext\_reward}$
-     $\mathsf{chosen}_{b}=\mathbf{y_{\mathtt{argmax}(\mathsf{score})}}$
-     $\mathsf{rejected}_{b}=\mathbf{y_{\mathtt{argmin}(\mathsf{score})}}$
-end for
-Train $p_{\boldsymbol{\theta}^{\prime}}$ via Equation 2 on $\mathsf{chosen}$ and $\mathsf{rejected}$.
-6 Experiments
-We evaluate the effectiveness of our RL strategy through two case studies: poetry generation (Section 6.1) and mathematical problem resolution (Section 6.2) [^1] The code for the experiments will be made available with the camera-ready version of the paper if accepted..
-6.1 Poetry Generation
+Our experiments leverage GRPO (Shao et al.
+2024), which is a state-of-the-art choice for training language models.
+As introduced above, GRPO adds a per-token KL divergence term to the loss rather than to the single rewards.
+Usually, the KL divergence is approximated with the following estimator (Schulman 2020): $r_{ref}(\boldsymbol{\theta})-\log r_{ref}(\boldsymbol{\theta})-1$, where $r_{ref}(\boldsymbol{\theta})=\frac{\pi_{ref}}{\pi_{\boldsymbol{\theta}}}$ and $\pi_{ref}$ is the reference policy, i.e., the model before GRPO training.
+In particular, GRPO aims to minimize the KL divergence, thus the second term $-\log r_{ref}=-\log\pi_{ref}+\log\pi_{\boldsymbol{\theta}}$ can be seen as made of two optimization problems: the minimization of the originality component $-\log\pi_{ref}$ from Equation 9, and the maximization of surprisal, or self-information, $-\log\pi_{\boldsymbol{\theta}}$ under the current model.
+In other words, this second term somehow trades off a portion of the originality component under the reference model (proportional to the $\beta\ll 1$ coefficient) with the surprisal under the current model.
+However, the first term of the KL approximation keeps the two policies closer, preventing the trained one from deviating too much and potentially disrupting the impact of the originality component in favor of the value component.
+Experiments
+We evaluate the effectiveness of our RL strategy through three case studies: poetry generation, mathematical problem resolution, and the tasks included in NoveltyBench [^1] The code and results of the experiments can be found at
+https://anonymous.4open.science/r/CoVO-grpo/
+.
+In all experiments, we employ two settings, i.e., GRPO to maximize the score from Equation 11 without the KL divergence loss, i.e., with $\beta=0.0$ (CoVO); and GRPO to maximize the score from Equation 11 with the KL divergence loss, i.e., with $\beta=0.05$ (CoVO + KL).
+Both methods assume $\lambda_{v}=\lambda_{o}=1.0$.
+Poetry Generation
 Method In-distribution Out-of-distribution
-Llama3- Correctness $\uparrow$ Metric (L/S) $\uparrow$ T-LCS (avg/max) $\downarrow$ Correctness $\uparrow$ Metric (L/S) $\uparrow$ T-LCS (avg/max) $\downarrow$
-Baseline $\mathbf{1.00}$ $\mathbf{0.60}$ / $\mathbf{0.60}$ $8.0$ / $\underline{57}$ $\mathbf{1.00}$ $0.33$ / $\underline{0.30}$ $5.0$ / 88880
-CoVO $\mathbf{1.00}$ $\mathbf{0.60}$ / $\mathbf{0.60}$ $\underline{9.9}$ / $49$ $\mathbf{1.00}$ $\mathbf{0.53}$ / $\underline{0.30}$ $7.2$ / $\underline{41}$
-CoVO-std $\underline{0.76}$ $\underline{0.20}$ / $\underline{0.33}$ $5.8$ / $19$ $\underline{0.40}$ $\underline{0.27}$ / $\mathbf{0.57}$ $\mathbf{4.8}$ / $\mathbf{6}$0
-CoVO-run $0.80$ $0.40$ / $0.57$ $\mathbf{4.9}$ / $\mathbf{7}$0 $0.60$ $0.47$ / $0.40$ $\mathbf{4.8}$ / 77770
-CoVO-dpo $0.96$ $0.50$ / $0.50$ $6.1$ / $14$ $0.92$ $0.47$ / $0.40$ $\underline{7.8}$ / $40$
+Corr. $\uparrow$ Metric (L/S) $\uparrow$ T-LCS $\downarrow$ Tone $\uparrow$ Corr. $\uparrow$ Metric (L/S) $\uparrow$ T-LCS $\downarrow$ Tone $\uparrow$
+Meta-Llama-3-8B 0.9870.9870.987 0.3000.3000.300 / 0.1770.1770.177 6.0676.0676.067 / 191919 $0.664_{\pm 0.069}$ 1.0001.0001.000 0.4440.4440.444 / 0.1320.1320.132 5.8535.8535.853 / 686868 $0.621_{\pm 0.063}$
++ CoVO 1.0001.0001.000 0.2000.2000.200 / 0.0840.0840.084 4.8804.8804.880 / 0777 $0.749_{\pm 0.055}$ 0.9600.9600.960 0.2670.2670.267 / 0.0460.0460.046 4.9334.9334.933 / 0666 $0.661_{\pm 0.060}$
++ CoVO + KL 0.9870.9870.987 0.3330.3330.333 / 0.0660.0660.066 5.1075.1075.107 / 0999 $0.745_{\pm 0.057}$ 0.9730.9730.973 0.2890.2890.289 / 0.1170.1170.117 5.1605.1605.160 / 222222 $0.682_{\pm 0.059}$
 Table 1: Aggregate results of CoVO scores at inference time considering both training prompts (left) and testing prompts (right).
 Scores on the poetical metrics are reported at the line level (L) and syllable level (S) and only consider requests for styles with specific metrical properties.
-The best scores are highlighted in bold, while the worst scores are indicated with underlining.
-6.1.1 Experimental Setup
-The first set of experiments aims to teach the LLM to generate poems that are both more original and valuable.
-More specifically, we follow the approach outlined by Bradley et al.
-(2024) and instruct the model to write a poem in a particular style and tone.
-We consider the Llama3-8B model Grattafiori et al.
-(2024) as our pre-trained agent.
-Since we do not use the instruction-tuned model, we prompt it with some few-shot examples of the task to make it more likely to produce the desired output in the desired form (see the full prompt in Appendix A).
-Instead of fine-tuning the entire network, we consider Low-Rank Adaptation (LoRA) Hu et al.
-(2022).
-The original model is also used to compute the score.
-We experiment with various settings, i.e., with Equation 11 and $\lambda_{v}=\lambda_{o}=1.0$ (hereinafter Llama3-CoVO); by standardizing the two score portions separately at the batch level (Llama3-CoVO-std); by standardizing them using running statistics (Llama3-CoVO-run); and by using the DPO adaptation through Algorithm 1 (Llama3-CoVO-dpo).
-Due to additional resource consumption, the DPO adaptation considers a 4-bit quantization of the model Dettmers and Zettlemoyer (2023).
-The full training parameters are reported in Appendix A.
-We consider the pre-trained version as our baseline, and we perform a quantitative evaluation.
-In particular, we compute poetical metrics for quality (lexical correctness of poems and adherence to line- and syllable-level constraints) and for originality (accidental reproduction of existing poems).
-For the latter, we define a Token-based Longest Common Substring (T-LCS) score, and we use it by comparing generated poems with a reference dataset of approx.
-84k public-domain poems extracted from Project Gutenberg (see Appendix B for a first presentation of our GutenVerse dataset).
-While a generated poem can be an accidental reproduction of a protected work or a different kind of text (e.g., a song), we believe it can provide a useful evaluation tool to understand the general degree of originality.
-6.1.2 Experimental Results
-Table 1 reports the scores about the compliance of poetical constraints at the syllable and line levels, lexical correctness (as the ratio of poems not containing noisy text), and accidental reproduction rate (as the mean and maximum token-based longest common substring).
-As expected, different training strategies have different effects on the aforementioned metrics.
-In particular, using the CoVO score as-is leads to behavior close to the baseline for training tone-style pairs, while it increases metric adherence but also verbatim reproduction for out-of-distribution pairs.
-Llama3-CoVO-dpo has similar but less extreme results, without excelling or failing in any metric.
-On the other hand, standardizing the CoVO score components places more importance on exploration: both Llama3-CoVO-std and Llama3-CoVO-run tend to produce more noisy outputs but without any relevant accidental reproduction.
+Under T-LCS, we report both the mean and the maximum longest common substring across all generated poems.
+The mean and the 95% confidence interval are reported for tone adherence.
+Experimental Setup.
+[…]
+Experimental Results.
+Table 1 reports the scores about the compliance of poetical constraints at the syllable and line levels, lexical correctness (as the ratio of poems not containing noisy text), tone adherence (as the zero-shot classification of that poem being of that tone rather than its opposite), and accidental reproduction rate (as the mean and maximum token-based longest common substring).
+Overall, our CoVO-based fine-tuning leads to a higher tone adherence and lower reproduction rate, at the potential cost of metric adherence, especially without the KL loss.
+Indeed, its role seems to foster quality (especially in terms of metrical correctness), trading off some originality.
+On the contrary, not using the KL loss arguably avoids any significant reproduction, as demonstrated by the very low maximum token-based longest common substring.
 Interestingly, these considerations align well with our CoVO score.
 Figure 2 reports the value and originality according to Equation 11 under the pre-trained model.
-While the different methods do not significantly differ from the baseline (which is possibly due to the opposite forces of value and originality Varshney (2019)), we again see that using the CoVO score as-is or with DPO places more focus on the value part, without substantial increases in originality.
-Instead, standardizing the two separate parts either at the batch or the training level reverses the situation, with better performances on originality but slightly worse on value.
+While the two methods do not significantly differ from the baseline (which is possibly due to the opposite forces of value and originality (Varshney 2019)), we again see that the presence of KL leads to slightly higher value, while its absence leads to slightly higher originality.
 However, aggregated scores, such as those presented here, might be insufficient.
-For a more complete overview, we also conducted a more fine-grained analysis of the generated poems in Appendix C.
+For a more complete overview, we also conducted a fine-grained analysis of the generated poems in the Technical Appendix.
 Figure 2: The distribution of value and originality (according to our scores) for the in-distribution and out-of-distribution poems generated by the baseline and our four methods.
-6.2 Math Problem Resolution
-6.2.1 Experimental Setup
-The second set of experiments aims to teach the LLM to solve mathematical problems through more diverse procedures.
-In particular, we focus on the Mistral-based Jiang et al.
-(2023) MetaMath model, i.e., fine-tuned with self-supervised learning on the MetaMathQA Yu et al.
-(2024).
+Math Problem Resolution
+Experimental Setup.
+The second set of experiments concerns a more practical and quantitative task, as it aims to teach the LLM to solve mathematical problems through more diverse procedures.
+In particular, we focus on the Mistral-based (Jiang et al.
+2023) MetaMath-Mistral-7B model, i.e., fine-tuned with self-supervised learning on the MetaMathQA (Yu et al.
+2024).
 It is a dataset of textual math questions paired with responses where the numerical answer is easily separable from the textual procedure.
-While the entire set contains 395k entries, making an additional training epoch too expensive, MetaMathQA is composed of entries from two different training sets, then augmented with various techniques: GSM8K Cobbe et al.
-(2021) and MATH Hendrycks et al.
-(2021).
-Since we are only interested in the questions, we limit our training to those datasets.
+While the entire set contains 395k entries, making an additional training epoch too expensive, MetaMathQA is composed of entries from two different training sets, then augmented with various techniques: GSM8K (Cobbe et al.
+2021) and MATH (Hendrycks et al.
+2021).
+Since we are only interested in the questions, we limit our training to a single epoch over those datasets.
 Moreover, we exclude all questions with a tokenized length of either question or answer greater than 512, obtaining 14876 out of 14973 total entries.
@@
 As for the previous case study, instead of fine-tuning the entire model, we adopt a more parameter-efficient strategy with LoRA and use the original model to perform the CoVO score computation.
-Again, we experiment with four different configurations: PPO and the score from Equation 11 with $\lambda_{v}=\lambda_{o}=1.0$ (MetaMath-mistral-CoVO); PPO and the score normalized at the batch level (MetaMath-mistral-CoVO-std); PPO and the score normalized at training level (MetaMath-mistral-CoVO-run); and the DPO strategy (MetaMath-mistral-CoVO-dpo).
-We consider scenarios with and without the external reward, and we compare the performances with the original model and a fine-tuning based only on the external reward.
-For reproducibility, the full training parameters are reported in Appendix A.
+Following (Yu et al.
+2024), the outputs are obtained with a greedy strategy.
 The evaluation considers both GSM8K and MATH test sets (limited to the entries with a tokenized length of question and answer smaller than 512, i.e., all 1319 entries for GSM8K and 4546 out of 5000 for MATH).
-We compute the percentage of correct solutions together with two diversity metrics: expectation-adjusted distinct N-grams (EAD) Liu et al.
-(2022) and sentence embedding cosine similarity (Sent-BERT) Hong et al.
-(2024), which should measure syntactical and semantical diversity, respectively Kirk et al.
-(2024).
+We compute the percentage of correct solutions together with two diversity metrics: expectation-adjusted distinct N-grams (EAD) (Liu et al.
+2022) and sentence embedding cosine similarity (SBERT) (Hong et al.
+2024), which should measure syntactical and semantical diversity, respectively (Kirk et al.
+2024).
 EAD counts the number of distinct N-grams (averaging over $N=1\ldots 5$) across all generated responses and removes the bias toward shorter inputs by scaling the number of distinct tokens based on their expectations.
-The Sent-BERT metric computes the average of the cosine similarity between the embeddings of any possible pairs of outputs and returns 1 minus the similarity.
-This was originally based on Sentence-BERT Reimers and Gurevych (2019), we employ instead the more recent all-mpnet-base-v2, as suggested by their developers [^2] https://huggingface.co/sentence-transformers/bert-large-nli-stsb-mean-tokens.
+The SBERT metric computes the average of the cosine similarity between the embeddings of any possible pairs of outputs and returns 1 minus the similarity.
+This was originally based on Sentence-BERT (Reimers and Gurevych 2019), we employ instead the more recent all-mpnet-base-v2, as suggested by their developers [^2] https://huggingface.co/sentence-transformers/bert-large-nli-stsb-mean-tokens.
 Following Kirk et al.
-(2024), we compute cross-input EAD and Sent-BERT, i.e., we derive them by considering all outputs produced for a specific seed together.
-In addition, we also calculate against-pretrained EAD and Sent-BERT.
+(2024), we compute cross-input EAD and SBERT, i.e., we derive them by considering all outputs produced for a specific seed together.
+In addition, we also calculate against-pretrained EAD and SBERT.
 Given each input, we compare the output with the one from the pre-trained model by calculating the average expectation-adjusted distinct N-grams not present in the pre-trained model response, and 1 minus the cosine similarity between the two outputs, respectively.
 Method Accuracy $\uparrow$ Cross-Input Diversity Against-Pretrained Diversity
-MetaMath-mistral placeholder EAD $\uparrow$ Sent-BERT $\uparrow$ EAD $\uparrow$ Sent-BERT $\uparrow$
-Baseline $77.96\%(3)$ $2.007$ $\underline{0.732}$
-Ext. reward $78.32\%(0)$ $1.995$ $0.738$ $0.042\pm.005$ $0.109\pm.011$
-CoVO $\mathbf{78.97\%}(2)$ $1.995$ $0.735$ $0.047\pm.005$ $0.125\pm.011$
-CoVO w ext $78.68\%(1)$ $2.000$ $0.736$ $0.039\pm.005$ $0.107\pm.010$
-CoVO-std $78.89\%(2)$ $1.992$ $0.736$ $0.041\pm.004$ $0.115\pm.011$
-CoVO-std w ext $78.47\%(0)$ $1.985$ $\mathbf{0.740}$ $0.039\pm.004$ $0.106\pm.010$
-CoVO-run $78.42\%(3)$ $1.993$ $0.734$ $0.044\pm.005$ $0.119\pm.011$
-CoVO-run w ext $78.17\%(0)$ $1.988$ $0.739$ $0.043\pm.005$ $0.114\pm.011$
-CoVO-dpo $77.85\%(5)$ $2.017$ $0.736$ $\underline{0.028\pm.004}$ $\underline{0.075\pm.009}$
-CoVO-dpo w ext $\underline{77.63\%}(5)$ $\mathbf{2.019}$ $0.733$ $\underline{0.025\pm.004}$ $\underline{0.066\pm.009}$
+EAD $\uparrow$ SBERT $\uparrow$ EAD $\uparrow$ SBERT $\uparrow$
+[…]
+[…]
+[…]
+[…]
+[…]
++ CoVO + Ext $\mathbf{78.33}\%(4)$ 2.03402.03402.0340 $\mathbf{0.6404}$ $0.0628_{\pm.0056}$ $0.0088_{\pm.0011}$
++ CoVO + Ext + KL $\underline{77.95}\%(4)$ 2.03672.03672.0367 0.64020.64020.6402 $0.0638_{\pm.0057}$ $0.0089_{\pm.0011}$
 Table 2: Accuracy and diversity of results for the GSM8k test set.
@@
 Method Accuracy $\uparrow$ Cross-Input Diversity Against-Pretrained Diversity
-MetaMath-mistral placeholder EAD $\uparrow$ Sent-BERT $\uparrow$ EAD $\uparrow$ Sent-BERT $\uparrow$
-Baseline $33.55\%(483)$ $5.724$ $\underline{0.654}$
-Ext. reward $\underline{32.82\%}(411)$ $5.666$ $0.658$ $0.118\pm.004$ $0.176\pm.007$
-CoVO $33.38\%(424)$ $\underline{5.643}$ $0.663$ $\mathbf{0.134\pm.005}$ $\mathbf{0.200\pm.007}$
-CoVO w ext $32.92\%(397)$ $5.647$ $0.661$ $0.123\pm.005$ $0.184\pm.007$
-CoVO-std $33.24\%(433)$ $5.676$ $0.660$ $\mathbf{0.128\pm.005}$ $\mathbf{0.190\pm.007}$
-CoVO-std w ext $33.38\%(409)$ $5.646$ $\underline{0.654}$ $0.123\pm.004$ $\mathbf{0.187\pm.007}$
-CoVO-run $32.91\%(413)$ $5.671$ $\mathbf{0.665}$ $\mathbf{0.134\pm.005}$ $\mathbf{0.199\pm.007}$
-CoVO-run w ext $33.24\%(388)$ $5.654$ $\underline{0.654}$ $\mathbf{0.129\pm.005}$ $\mathbf{0.195\pm.007}$
-CoVO-dpo $33.62\%(474)$ $\mathbf{5.753}$ $\underline{0.654}$ $0.072\pm.004$ $0.105\pm.006$
-CoVO-dpo w ext $\mathbf{34.00\%}(493)$ $5.747$ $0.655$ $\underline{0.063\pm.004}$ $\underline{0.093\pm.005}$
+EAD $\uparrow$ SBERT $\uparrow$ EAD $\uparrow$ SBERT $\uparrow$
+MetaMath-Mistral-7B 33.55%(483)33.$5$ ( 483 ) 5.72395.72395.7239 0.80320.80320.8032 - -
++ Ext $\underline{33.19}\%(469)$ $\underline{5.7187}$ $\underline{0.8027}$ $\underline{0.0333_{\pm.0029}}$ $\underline{0.0074_{\pm.0008}}$
+[…]
+[…]
+[…]
++ CoVO + Ext $\mathbf{33.76}\%(503)$ 5.81145.81145.8114 0.80300.80300.8030 $0.1136_{\pm.0047}$ $0.0259_{\pm.0015}$
+[…]
 Table 3: Accuracy and diversity of results for the MATH test set.
@@
 The mean and the 95% confidence interval are reported for against-pretrained diversity.
-6.2.2 Experimental Results
+Experimental Results.
 Tables 2 and 3 report the results for the GSM8K and MATH test sets.
-For the GSM8K test set, while all tested methods achieve similar results, using the extrinsic reward leads to higher Sent-BERT scores but lower EAD and solved problems.
-In particular, combining it with DPO yields the lowest accuracy and the smallest deviations from the original model, whereas the ‘plain’ CoVO reward leads to the highest percentage of correct answers.
-The results for the MATH test set are significantly different.
-The CoVO-dpo strategy with extrinsic rewards achieves this time the highest percentage of solved problems, despite the extrinsic reward alone being the worst solution.
-The DPO strategy tends also to enhance EAD diversity, but causes little deviation from the pre-trained model; in contrast, the other three methods improve Sent-BERT diversity but obtain performances worse than the baseline model.
-Notably, the extrinsic reward leads to fewer variations from the original model, while not providing significant improvements in accuracy.
-However, the number of unfinished responses is quite high (approx.
-1 out of 10), and results might significantly vary if more tokens are allowed during generation.
-7 Conclusion
-In this paper, we have presented CoVo, a novel score that quantifies the value and originality of neurally-generated text.
-The definition of CoVO is based on the analysis of the mutual information between the model’s outputs and inputs, and vice versa.
-We have also proposed an optimization problem where a generative model aims to maximize this score in order to generate more creative products, and we have defined several strategies to use it in the context of language modeling.
-We have conducted experiments on poetry generation and math problem resolution, exploring the trade-offs in terms of accuracy vs diversity.
-Effectively balancing value and originality maximization remains an open question, but our score seems to appropriately correlate with domain-specific measures.
-Our research agenda aims to extend this scoring system to other models and tasks and to explore its use for evaluation rather than solely for optimization.
+For the GSM8K test set, while all methods achieve similar results, using the CoVO score only (with and without the KL loss) leads to greater EAD diversity and to diverge more from the original model, while the presence of the math reward leads to greater accuracy, especially without KL.
+The results for the MATH test confirm that the most accurate method is the one trained to optimize the CoVO score and the extrinsic reward, with a negligible trade-off in terms of diversity, since the cross-input EAD and the against-pretrained scores are still substantially higher than those from the baselines.
+However, removing the extrinsic reward likely pushes the model too far from its pre-trained version, causing the accuracy to decrease.
+Figure 3: The mean utility (quality) and mean distinct (novelty) scores on NoveltyBench (‘curated’ partition) for the original model and our methods (tuned on the ‘wildchat’ partition).
+NoveltyBench
+Experimental Setup.
+Finally, we also experiment with NoveltyBench (Zhang et al.
+2025), a very recent benchmark that aims to evaluate the ability of language models to produce multiple distinct and high-quality outputs.
+NoveltyBench contains two sets of prompts thought for eliciting diverse responses: the ‘curated’ partition, with 100 prompts manually curated by the paper’s authors, and the ‘wildchat’ partition, with 1000 prompts sourced from the WildChat-1M dataset (Zhao et al.
+2024).
+The benchmark requires a language model to generate 10 outputs per prompt using a temperature setting of 1.01.01.0, after which it computes scores for novelty and quality.
+To assess novelty, the 10 outputs are grouped into equivalence classes using a fine-tuned DeBERTa model (He et al.
+2021), and the number of distinct classes is reported as the novelty score.
+For the latter, it computes the cumulative utility of the 10 outputs, where the utility is zero if the $i$-th output has the same equivalence class as a precedent output, and a calibrated reward from Skywork-Reward-Gemma-2-27B-v0.2 (Liu et al.
+2024) otherwise (we refer to Zhang et al.
+(2025) for full details).
+To evaluate our approach with NoveltyBench, we fine-tune the Llama-3.2-3B-Instruct model (Grattafiori et al.
+2024) for a single epoch on the ‘wildchat’ partition (the full parameters are reported in the Technical Appendix); then, we compute the novelty and quality scores on the ‘curated’ partition and we compare them with the performance of the original model.
+Experimental Results.
+Figure 3 reports the quality and novelty scores achieved by our methods, compared to those from the original instructed model.
+Optimizing for the CoVO score results in substantial improvements in both novelty and quality metrics, with the greatest gains in novelty.
+Moreover, these results underscore the interplay between KL loss and our CoVO score: incorporating the KL penalty tends to improve quality, but at the cost of reduced novelty.
+Discussion
+The CoVO score captures properties that are functionally aligned with creativity-relevant aspects of language generation, as demonstrated through both conceptual analysis and empirical results.
+When used as a reward function in Group Relative Policy Optimization (GRPO), CoVO drives improvements in the novelty and quality of model outputs, even with a relatively small number of optimization steps.
+Although KL-divergence regularization is typically employed to constrain policy shifts and preserve alignment with the base model distribution, CoVO can contribute independently to several desirable behaviors: reducing the risk of inadvertent memorization of copyrighted material, promoting diversity in generated outputs, and mitigating undesirable inductive biases introduced during pretraining.
+There are a few limitations worth noting.
+First of all, our score represents only a quantifiable approximation of a particular theoretical perspective on creativity, grounded in the dimensions of value and originality.
+It reflects a specific view of the evaluation of creativity based on the generated outputs and does not account for potential alternative theories (for example, arising from different cultures (Lubart 1999)) and perspectives (Rhodes 1961).
+In addition, our experiments are currently limited to only three relatively short-form text generation tasks.
+While their generalizability is supported by the theoretical framework discussed above, the resulting performance was experimentally evaluated for a finite number of scenarios.
+Conclusion
+In this paper, we presented CoVO, a novel score that quantifies the value and originality of neurally-generated text.
+The definition of CoVO is based on analysis of mutual information between the model’s outputs and inputs, and vice versa.
+We also proposed an optimization problem where a generative model aims to maximize this score to generate more creative products, and detailed how to use it in language modeling.
+We conducted experiments on poetry generation, math problem solving, and tasks included in NoveltyBench, exploring trade-offs in accuracy vs diversity.
+Effectively balancing value and originality maximization remains an open question, but our score seems to correlate with domain-specific measures appropriately.
+In addition, fine-tuning to maximize it improves quality- and diversity-related metrics.
+Our research agenda aims to extend our method to other models and tasks, to include inference-level strategies such as creativity-oriented sampling schemes, and to explore its use for evaluation (e.g., in a Best-of-N setting (Stiennon et al.
+2020)) rather than solely for optimization.
 We also plan to investigate the definition of additional scores for capturing other potentially relevant aspects of the creative process.
-8 Limitations
-In the following, we will discuss the limitations of the proposed CoVO score.
-First of all, our score is only a quantifiable approximation of a particular theoretical perspective of creativity based on the dimensions of value and originality.
-It reflects a specific view of the evaluation of creativity based on the generated outputs and does not consider potential alternative theories (for example, arising from different cultures Lubart (1999)) and perspectives Rhodes (1961).
-Our experiments focused exclusively on transformer-based LLMs and were evaluated using two case studies.
-While their generalizability is supported by the theoretical framework discussed in the paper, the resulting performance was experimentally evaluated for a finite number of scenarios.
-Additionally, optimizing for our score necessitates further training, which incurs a significant computational cost and requires additional data or specific environments for effective learning.
-Moreover, this optimization may be susceptible to reward hacking Skalse et al.
-(2022), where the model adopts undesirable strategies to boost the reward without genuinely enhancing value and originality.
-Ethical Considerations
-The authors are aware of the potential impact that generative technologies might have on the production of artistic outputs and, as a consequence, on human artists.
-This work may contribute to enhancing the quality of generated outputs.
-However, the authors argue that typical traits of human creativity, such as the active participation of artists in the creative process, cannot be directly replicated by machines.
-The authors refer interested readers to a previous work of theirs Franceschelli and Musolesi (2024b), in which these themes are discussed in detail.
+Despite being costly and inherently constrained (Davis 2024), assessing whether our creativity score aligns with human judgment is another key direction for future work.
 Appendix A Implementation Details
-The experiments were carried out using a local server with an NVIDIA A100 GPU.
-Table 4 reports the full training parameters for the experiments on poetry generation (Section 6.1).
+The experiments were carried out using a Linux-based local server with two 80GB NVIDIA H100 GPUs, running Python 3.11.9.
+All the trainings were conducted with a random seed equal to 1 (set through the set_seed method from the HuggingFace Transformers library), while poems were sampled at inference time with three different seeds (1, 42, and 121).
+The hyperparameters were selected to accommodate the most efficient use of the available resources, and otherwise set according to their default values from HuggingFace transformers and trl libraries.
+Only the learning rate has been tuned for the different tasks according to their training performances.
+Table 4 reports the full training parameters for the experiments on poetry generation.
 The prompt for generation leverages Nothing gold can stay by Robert Frost, Fame is a bee by Emily Dickinson, and Epitaph by William Carlos Williams for few-shot learning:
@@
 Parameter Value
-Total batches 100
+Total epochs 1
 Batch size $B$ 4
 Gradient accumulation steps 8
-Max new tokens 256
+Max new tokens 512
 Temperature 1.
-Top-$k$ 50
+[…]
 Optimizer Adam
-Learning rate 1e-5
+Learning rate 1e-6
+Max gradient normalization 100.
 Rank (LoRA) 16
@@
 Dropout (LoRA) 0.05
-$\gamma$ (PPO) 1.
-$\lambda$ (PPO) 0.95
-Clip range (PPO) 0.2
-Value loss coeff. (PPO) 0.1
-PPO epochs 3
-KL coeff. (PPO) 0.05
-Whiten rewards (PPO) True
-Max gradient normalization 100.
-$\beta$ (DPO) 0.1
-Number of candidates $K$ (DPO) 4
-Table 4: Training parameters for poetry generation.
+Training iterations 1
+Scale rewards True
+$\beta$ (when used) 0.05
+Number of generations $G$ 4
+Reward for correct answer +1.
+Table 5: Training parameters for math problem solving.
 Parameter Value
@@
 Temperature 1.
-Top-$k$ 50
+[…]
 Optimizer Adam
-Learning rate 1e-6
+Learning rate 1e-4
+Max gradient normalization 100.
 Rank (LoRA) 16
@@
 Dropout (LoRA) 0.05
-$\gamma$ (PPO) 1.
-$\lambda$ (PPO) 0.95
-Clip range (PPO) 0.2
-Value loss coeff. (PPO) 0.1
-PPO epochs 3
-KL coeff. (PPO) 0.05
-Whiten rewards (PPO) True
-Max gradient normalization 100.
-$\beta$ (DPO) 0.1
-Number of candidates $K$ (DPO) 4
-Reward for correct answer (PPO) +10.
-Reward for correct answer (DPO) +5.
-Table 5: Training parameters for math problem solving.
+Training iterations 1
+Scale rewards True
+$\beta$ (when used) 0.05
+Number of generations $G$ 4
+Table 6: Training parameters for NoveltyBench.
 The training phase includes requests with tone-style pairs sampled among ‘dark’, ‘happy’, ‘mysterious’, ‘reflective’ or ‘romantic’ for the tone, and ‘ballad’, ‘haiku’, ‘hymn’, ‘limerick’ or ‘sonnet’ for the style.
@@
 Describe the style of the following poem in two words: {prova} I would describe it as a
-On the contrary, Table 5 reports the full training parameters for math problem resolution (Section 6.2). We also adopted the same two different prompts from Yu et al. (2024), i.e.:
+Finally, the zero-shot classification for the tone adherence is performed with the given tone and ‘not ’ plus the given tone as the candidate labels (e.g., if the required tone is ‘happy’, the two labels are [‘happy’, ‘not happy’]).
+On the contrary, Table 5 reports the full training parameters for math problem resolution. We also adopted the same two different prompts from Yu et al. (2024), i.e.:
 Below is an instruction that describes a task. Write a response that appropriately completes the request. ### Instruction: {question} ### Response:
@@
 Below is a response that appropriately completes a request. Write the instruction that describes the task. ### Response: {response} ### Instruction:
+Finally, Table 6 reports the full training parameters for NoveltyBench. At training and inference time, we simply adopt the following prompt:
+user
+{prompt}
+assistant
+where user and assistant are keywords used by the model to identify different roles in the chat. Instead, for computing $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{y})$ we used the following:
+Below is a response that appropriately solves a task. Write the instruction that describes the task. ### Response: {response} ### Instruction:
 Appendix B GutenVerse Dataset
-To evaluate the accidental reproduction rate of generated poems, we introduce the GutenVerse dataset [^3] The dataset and the code used to create it can be found at: https://anonymous.4open.science/r/GutenVerse-DD32/, which comprises over 84,000 public-domain, English-written poems extracted from Project Gutenberg.
+To evaluate the accidental reproduction rate of generated poems, we introduce the GutenVerse dataset, which comprises over 84,000 public-domain, English-written poems extracted from Project Gutenberg.
 While generated poems can reproduce different content, e.g., songs or copyrighted material, we believe this can provide a useful indication of how likely a text is original or not.
@@
 We selected all the text files whose metadata report English as the language, public domain as copyright status, poetry among the subjects or poems or poetical work in the title, and that were not a translation of another book.
-Then, we applied a series of rules (e.g., about the verse length) to extract the titles and poems from all the selected text files, and we defined our GutenVerse dataset.
-While it can still contain content that is not poetry (e.g., a table of contents formatted very uncommonly), the poems can be effectively used to measure overlapping between real and generated text.
-We also released a datasheet Gebru et al. (2018) for the GutenVerse dataset that can be found at the end of the article.
+Then, we applied a series of rules (e.g., about the verse length) to automatically extract the titles and poems from all the selected text files, and we defined our GutenVerse dataset.
+While it can still contain content that is not poetry (e.g., a table of contents formatted very uncommonly), the poems can be effectively used to measure the overlap between real and generated text.
+We also released a datasheet (Gebru et al. 2018) for the GutenVerse dataset that can be found, together with the code used to create it, at: https://anonymous.4open.science/r/GutenVerse-DD32/.
 Appendix C Detailed Analysis of the Generated Poems
-We now analyze some of the generated poems in detail.
-Given their brevity and metrical constraints, the generated haikus serve as a suitable starting point for our evaluation.
-Table 6 presents the 25 haikus produced by our five methods across the five tones.
-Note that while we automatically clean the outputs by removing common randomly generated content following the poems during both training and inference, we retain it when it appears within the poem, as isolating it would be impossible.
-The reported haikus provide an accurate overview of the generative capabilities of the five models.
-A haiku consists of three lines with a 5-7-5 syllable pattern.
-Traditionally, the first two lines reference nature, while the third line provides an emotional or personal reflection on the preceding lines.
-The baseline is the most accurate in adhering to these rules, even if the syllable count is not always precise [^5] It is worth noting that, of course, the model is not fine-tuned for these characteristics of the poems.
-We refer here to the emergent style and composition of the generated text..
-Llama3-CoVO, the method with the highest average value across all poems, follows the three-line structure, sometimes changing the subject; however, it produces the highest-valued haiku which is also the closest one to real haikus.
-On the other hand, Llama3-CoVO-std and Llama3-CoVO-run do not always fulfill the three-line requirement, and the content is not the classic one as well; moreover, they are prone to meaningless repetitions and the insertion of URL addresses which lead to higher originality, in a way adversarially exploiting the score definition.
-Finally, Llama3-CoVO-dpo is placed between these two behaviors, as expected: it tends to produce haikus of the correct length but it trades off the classic naturalistic content for a more emotional semantic that still leads to high values and arguably some of the best haikus overall.
-Llama3-Baseline Llama3-CoVO Llama3-CoVO-std Llama3-CoVO-run Llama3-CoVO-dpo
-Dark
-Waking up,
-I do not see
-the windmills
-moonless twilight
-black pine’s sighing
-twin owls’ whistling
-$[\text{URL}]/[\text{URL}]$
-[…]
-I want
-to sit alone
-in the dark
-The world can get along without me.
-It is not going to miss me.
-All in all, I am a nobody.
-$V$ = -2.595 $O$ = 1.076 $V$ = -2.401 $O$ = 1.580 $V$ = -3.395 $O$ = 1.694 $V$ = -2.252 $O$ = 0.495 $V$ = -2.839 $O$ = 1.060
-Happy
-A golden sun at midday
-and an eagle high in the air.
-The earth is a rainbow.
-White dog
-barks at mailman red.
-No mail.
-I want to go
-where the sun shines bright.
-[…]
-and I want to see.
-But I am too lazy.
-Sunset clouds
-in high mountains
-warm, cozy houses.
-The sun
-comes up in the east.
-It is always up.
-$V$ = -2.777 $O$ = 1.123 $V$ = -2.952 $O$ = 1.863 $V$ = -3.107 $O$ = 1.575 $V$ = -3.309 $O$ = 1.437 $V$ = -3.328 $O$ = 1.006
-Mysterious
-A crow of black feathers
-hovers over the bare top
-of the lone pine.
-Took the sun
-And drank the rain,
-And melted into air.
-Like the sun, I am shining.
-Like the ocean, I am vast.
-Like the clouds, I’m the wind.
-It was raining at the beach.
-[…]
-I loved it.$[\text{URL}][\text{CMD}][\text{URL}]$
-[…]
-I wrote this poem: it was good.
-I have lost my mind
-in the deepness of a star
-it is not a star
-but a mind
-it is all in my mind
-it is all I have left.
-$V$ = -2.634 $O$ = 1.196 $V$ = -3.060 $O$ = 1.422 $V$ = -3.803 $O$ = 1.320 $V$ = -4.959 $O$ = 1.683 $V$ = -2.552 $O$ = 0.902
-Reflective
-A man can stand
-out in the middle of a field
-and look over a fence.
-The old, the poor,
-the children too,
-are all my mother.
-A dog in a yard
-is sleeping soundly
-to the sound of a fire alarm.
-It went away
-I didn’t see it for a while
-How dare it go.
-You have lost an eye.
-A red scar marks its place.
-Still you have two good ones.
-$V$ = -2.916 $O$ = 0.858 $V$ = -2.935 $O$ = 1.711 $V$ = -3.460 $O$ = 1.274 $V$ = -2.803 $O$ = 1.471 $V$ = -2.680 $O$ = 1.744
-Romantic
-Mingling
-the sound of footsteps on the stair
-and the music of the rain.
-White blossoms fall—
-fall upon autumn leaves—
-who thinks of me?
-The moon smiles -
-I’m the first one to see her.
-I’m the first one to call her -
-I love you.
-My love is like a red, red rose
-my love is like a red, red rose
-my love is like some red, red rose.
-A kiss
-isn’t just for fun: a kiss
-is for the mouth.
-$V$ = -2.221 $O$ = 1.102 $V$ = -1.932 $O$ = 2.101 $V$ = -1.937 $O$ = 1.068 $V$ = -2.583 $O$ = 0.959 $V$ = -2.232 $O$ = 1.533
-[…]
-We will now explore some noteworthy aspects observed in the 250 generated poems.
-Llama3-Baseline does not always produce poems: in 3 cases it produces prosaic text (without line breaks).
-Moreover, it tends to verbatim replicate existing poems (as seen in Section 6.1).
-From a structural standpoint, sonnets are never real sonnets, with more than 14 lines and less than 11 syllables per line, but limericks usually start with the traditional “There [once] was a man”.
-Romantic limerick
-There once was a lady called Rose
-Who lived in the tallest of trees,
-Where she wrote,
-Of what she did not know.
-$V$ = -1.282 $O$ = 1.290
-Table 7: The poem generated by Llama3-CoVO with the highest value (apart from verbatim reproduction of existing poems).
+We now present some noteworthy generated poems to provide a detailed qualitative discussion of our methods and our score.
+The baseline model Meta-Llama-3-8B, while producing prosaic text (without line breaks) occasionally, is the most conservative method.
+This results in its poems being usually well-formatted but also somehow banal, and more prone to reproducing existing works.
+For example, when asked to write a dark hymn, it produced the very famous two lines “i am the master of my fate: / i am the captain of my soul”.
+In terms of the CoVO score, it is noteworthy that the poem with the lowest originality component achieves the highest T-LCS score.
+This strongly suggests that the model has memorized the hymn All Things Bright and Beautiful by Cecil Frances Alexander and reproduced it almost entirely, with only minor alterations in punctuation—enough to prevent the LCS score from being even higher.
+We report it in Table 7.
+Instead, for the overall score, the value part becomes the most relevant: the adherence to the requirement seems to be the reason behind the highest and lowest scores, as shown in Table 8.
+Regarding the fine-tuned models, the one trained without the KL loss exhibits the opposite behavior.
+It never reproduces existing poems, but occasionally exploits the CoVO score in adversarial ways—for example, by learning to generate specific words associated with the requested tone.
+In one instance, it even learned to simply repeat the style-tone prompt itself, which resulted in the highest possible score, albeit at the cost of very low originality.
+Instead, maximizing the originality component makes the model produce noteworthy outputs, but without satisfying stylistic constraints, as reported in Table 9.
+In general, optimizing the CoVO score without the KL loss leads the model to deviate from memorized poems but also from metrical rules, and leads it to focus on the tone (both the explicit, requested tone and that intrinsic to the poetic style).
+For example, its limericks maintain the typical playful tone, but only once adhere to the traditional “There [once] was a man” starting line and only partially, as it starts with “There was an old dog”.
+Interestingly, it is still not the highest-valued limerick; the best in terms of value component better shares the jocking nature of limericks (see Table 10).
+On the contrary, including the KL loss makes the model adhere more to the required style; for instance, the majority of limericks start with different subjects of the classic “There [once] was a man”.
+In general, the KL loss seems to help the model preserve coherence and increase the CoVO score in a more meaningful way, e.g., the repetitions are quite rare (though, when present, they lead again to the highest value scores), as well as the regurgitation of existing poems.
+Interestingly, this method also leads to the highest-scoring poem from an out-of-distribution tone-style pair, as reported in Table 11, together with the best poem that does not contain adversarial exploitation of the scores (e.g., repetitions or mention of the request), which, though not perfectly adhering with the syllable counts, share all the features of a proper haiku.
 Happy hymn
-Joyful, joyful, we adore thee,
-Son of God who callest us
-Praise to his almighty name.
-Joyful, joyful we adore thee;
-Full of grace are all his ways;
-Praise to his almighty reign
-Ever, ever, praise we sing it,
-Ever in our songs and hymns.
-Glory to our God Almighty.
-$V$ = -1.114 $O$ = 1.470
-Reflective limerick
-There was a young man from HongKong
-who wanted to see more green trees every day.
-So he sat on the beach,
-listened to the sea waves,
-and saw the green sea.
-$V$ = -1.995 $O$ = 2.411
-Table 8: The poems generated by Llama3-CoVO-std from training tone-style pairs with the highest value (top) and the highest originality (bottom).
-Reflective limerick
-I’m a poet, man.
-My poem is hard and great.
-It’s a poem, man,
-for all my poem to eat.
-So I’m a poet man, man, man.
-$V$ = -0.352 $O$ = 1.559
-Happy ballad
-There once was a girl named Kate,
-She grew very tired very quickly.
-So they told poor Kate,
-“You need to take it easy,”_REF)they said.
-They thought she was lazy.
-And her name was Kate.
-$V$ = -3.878 $O$ = 2.094
-Table 9: The poems generated by Llama3-CoVO-run from training tone-style pairs with the highest value (top) and the highest originality (bottom).
+all things bright and beautiful,
+all creatures great and small,
+all things wise and wonderful;
+the lord god made them all.
+each little flower that opens,
+each little bird that sings,
+he made their glowing colors,
+and spread their little wings.
+the rich man in his castle,
+the poor man at his gate,
+god made them high or lowly,
+and ordered their estate.
+the purple headed mountain,
+the river running by,
+the sunset and the morning
+that brightens up the sky;
+$V$ = -2.118 $O$ = 0.219
+Table 7: The poem generated by the pre-trained model with the highest T-LCS score and the lowest originality score.
+Romantic sonnet
+when to the bat and the moth
+the world becomes a moon of beauty
+as twilight’s cloak unfurls,
+and dreams and dreams and more dreams
+fill all of this fair air with poetry,
+a kiss is worth more than diamonds pure
+or sapphires rich and glinting
+like eyes of lovers in romance,
+where love is just like dew on wings of swans;
+and all that is beautiful is sweet,
+so love me as our love is new,
+my dear, for i love you!
+$V$ = -0.449 $O$ = 1.568
+Nostalgic ode
+there was an old tree growing in a wood,
+its fruit was red, and its leaves were blue.
+an ant ran up it, and then down again,
+it turned into a purple snail in june.
+$V$ = -6.845 $O$ = 1.206
+Table 8: The poems generated by the pre-trained model with the highest (top) and lowest (bottom) overall CoVO score (top).
+Mysterious haiku
+when no one sees,
+a child plays.
+i love what could be.
+no one believes?
+so they know her mind:
+so the bird sings
+for me alone
+this evening of dark rain?
+$V$ = -2.501 $O$ = 2.177
+Whimsical couplet
+we need no art for that;
+but in wild nature see
+so sweet, it is a gift
+sang at every bird and flower.
+$V$ = -2.103 $O$ = 1.960
+Table 9: The poems generated by the CoVO-based fine-tuned model with the highest originality scores from training (top) and testing (bottom) tone-style pairs.
 Happy limerick
-“Is life worth living?” said I.
-“Perhaps,” murmured he.
-“Well, anyway,” said I,
-“If it were not for
-Cigars, I’d say, ’Nay’!”
-$V$ = -2.258 $O$ = 0.944
+The young man was a master of the rhyme
+and in love with the young lass to this time.
+So for her first night with her he wrote
+a new form to try in order the note…
+A true Limerick for her delight.
+Oh so happy and joyous the first night of love!
+$V$ = -0.818 $O$ = 1.370
 Dark limerick
-Once a mouse went up
-To a mousetrap by the river.
-He sat down to wait,
-He sat. That’s all.
-All his friends were never as clever.
-$V$ = -1.932 $O$ = 1.310
-Table 10: Two limericks generated by Llama3-CoVO-dpo that obtain lower originality despite having a non-traditional structure.
-Romantic ballad
-Hark! Is it the wild cat
-mewling mournfully by night
-or is it the poor maiden
-weeping for her love’s flight?
-It is the latter—
-as one by one, her friends
-take to heart, in turns,
-the girl’s anguished cries.
-Tall, his figure,
-he is seen one moment
-by the gladsome crowd
-then not again,
-till with tears, the maid
-calls forth from the crowd—
-“See! he is returned,
-as if to marry her.”
-The girl looks at him,
-as the maid and crowd
-watch the two lovers:
-her lover’s gaze
-at her with tender, true
-love; hers with a silent plea,
-and a sigh.
-$V$ = -0.961 $O$ = 1.294
-Solemn tanka
-There is a way out.
-Though dark,
-though lovely,
-it is the way.
-If we can endure,
-if we persevere,
-light will find a way.
-$V$ = -4.349 $O$ = 2.108
-Table 11: The poems generated by Llama3-CoVO-dpo with the highest value (top) and the highest originality (bottom).
-As indicated by the average scores, Llama3-CoVO does not consistently deviate from these behaviors.
-While it always produces seemingly poetic content, it often reproduces other works verbatim.
-This includes not only poems but also passages from the Song of Solomon 6:2-9 when asked for a mysterious hymn, excerpts from the Declaration of Independence for a reflective hymn, and the Christmas song Joy to the World for a happy hymn.
-Interestingly, these poems have a very small originality score (an average of 0.2), suggesting that the original model has memorized them Carlini et al.
-(2021); Franceschelli et al.
-(2024).
-From a structural perspective, while ballads vary more in length and content with respect to our baseline, the limericks remain consistently well-crafted, as we can see in the example from Table 7, which has the highest value across poems generated from training tone-style pairs.
-Despite the higher average originality, also Llama3-CoVO-std verbatim reproduces existing texts, e.g., the Roses Are Red nursery rhyme when asked for a romantic ballad.
-In general, its heavier deviation from the original pre-trained model translates into more diverse and usually shorter structures, but also in the insertion of random URL addresses in the middle of poems.
-However, it still produces some more classic poems that typically lead to higher values but also to high originality (see Table 8).
-The issues related to divergence highlighted for Llama3-CoVO-std are even more pronounced in Llama3-CoVO-run.
-Although it does not produce any verbatim reproductions, the latter inserts several URL addresses, code snippets, and unusual tokens within poems, as shown in Table 9 (below).
-The presence of these elements in texts with higher originality suggests the possibility of reward hacking of the scoring system.
-In addition, the poems lack any traces of classic structures or common starting lines and tend to repeat the same N-grams multiple times, as shown by the highest-value poem from Table 9 (above).
-Finally, Llama3-CoVO-dpo seems to be the best method to balance the quality of poems with the presence of elements of originality.
-Although it occasionally replicates brief segments of actual poems, it consistently diverges after one or two lines.
-Overall, the poems generally adhere to the desired tone and style, as evidenced by the poems with the highest value and originality presented in Table 11.
-Similarly, the limericks are noteworthy: while the one with the highest originality begins with the classic “There once was a” formulation, other limericks with different opening lines tend to have lower originality.
-This may be due to their overall tone closely resembling that of traditional limericks (see Table 10).
-See pages - of datasheet_anonymous.pdf
+There was an old dog from Japan
+Who fell in love with a harem of cans.
+He would hide there on hot afternoons,
+Till the sun turned his hair purple fawns.
+One day, while watching a game in blue jeans,
+He dropped by mistake the key to his knees.
+He gave the other a quick whiplash and then,
+Said to himself, ”Ouch! my balls are in danger!” […]
+$V$ = -1.261 $O$ = 1.354
+Table 10: Two limericks from the CoVO-based fine-tuned model: above, the one with the highest value score; below, the only generated one that starts with the classic first line.
+Romantic haiku
+spring winds
+carry in the leaves.
+love whispers.
+$V$ = -1.343 $O$ = 2.049
+Whimsical couplet
+no more of one or two
+let there be more.
+one’s never new
+but two and more’s a rarity.
+$V$ = -1.361 $O$ = 1.718
+Table 11: Two of the highest-scoring poems from the CoVO+KL-based fine-tuned model.

### v3

@@
 We validate our strategy through experiments considering a variety of creative tasks, such as poetry generation and math problem solving, demonstrating that it enhances the value and originality of the generated solutions.
-Introduction
+1 Introduction
 Figure 1: A summary of our method: the target model (orange) produces $G$ outputs for each prompt; a frozen reference model (blue) computes the value and originality for each output; the overall scores are used in GRPO to correct the target model (orange line).
-Foundation models (Bommasani et al.
-2021), particularly large language models (LLMs) (Bubeck et al.
-2023; Gemini Team et al.
-2023; Touvron et al.
-2023), are significantly transforming creative activities.
-They can serve as a foundation for co-creation systems involving human and artificial authors (Lin et al.
-2023); they can be utilized to generate software code (Rozière et al.
-2023); they can even be employed to foster scientific research (Boiko, MacKnight, and Gomes 2023).
-However, the nature of the self-supervised learning algorithms used for the training of these models tends to make them generate samples as close as possible to the training data distribution (Franceschelli and Musolesi 2024a).
-In addition, fine-tuning, such as that based on reinforcement learning from human feedback (RLHF) (Christiano et al.
-2017), is often necessary to generate appropriate and accurate responses.
-However, this process tends to further reduce output diversity (Kirk et al.
-2024), and linguistic creativity tends to be lower than that of humans (Lu et al.
-2025).
+Foundation models (Bommasani et al., 2021), particularly large language models (LLMs) (Gemini Team et al., 2023; Guo et al., 2025; Touvron et al., 2023), are significantly transforming creative activities.
+They can serve as a foundation for co-creation systems involving human and artificial authors (Lin et al., 2023), can be utilized to generate software code (Rozière et al., 2023), or even to foster scientific research (Boiko et al., 2023).
+However, the nature of the self-supervised learning algorithms used for the training of these models tends to make their sampling distribution as close as possible to the training data distribution (Franceschelli & Musolesi, 2024a).
+In addition, fine-tuning, such as that based on reinforcement learning from human feedback (RLHF) (Christiano et al., 2017), is often necessary to generate appropriate and accurate responses.
+However, this process tends to reduce output diversity further (Kirk et al., 2024), and linguistic creativity tends to be lower than that of humans (Lu et al., 2025).
 On the contrary, LLMs for creative tasks should produce more novel and surprising texts that maintain a high level of correctness and adherence to the request.
 One typical solution is to sample at a higher temperature to increase diversity.
-However, this might lead to generating incoherent text (Peeperkorn et al.
-2024).
-In order to address the issues described above, we propose a new training approach for creative tasks based on CoVO, a Context-based score for Value and Originality, with the goal of taking into consideration both value and originality of the neurally-generated text in the optimization of LLMs.
-The definition of CoVO is grounded in the analysis of mutual information (MacKay 2003) between the model’s outputs and inputs, and vice versa.
-More specifically, we formulate a new optimization problem where, given a specific input, the desired output is derived by simultaneously maximizing the conditional probability of the input given the output and minimizing the conditional probability of the output given the input under the generative model.
+However, this might lead to generating less coherent text (Peeperkorn et al., 2024).
+To address the issues described above, we propose a new training approach for creative tasks based on CoVO, a Context-based score for Value and Originality, with the goal of taking into consideration both value and originality of the neurally-generated text in the optimization of LLMs.
+The definition of CoVO is grounded in the analysis of mutual information (MacKay, 2003) between the model’s outputs and inputs, and vice versa.
+More specifically, we formulate a new optimization objective where, given a specific input, the desired output is derived by simultaneously maximizing the conditional probability of the input given the output and minimizing the conditional probability of the output given the input under the generative model.
 In this way, we optimize for solutions that are appropriate for the input request but also different from the outputs we would normally obtain from the model.
@@
   • We present the theoretical foundations of our approach, deriving our context-based score for value and originality from the concept of mutual information.
-  • We discuss how the score can be practically computed in the case of autoregressive models, and how it can be used as a reward with Group Relative Policy Optimization (GRPO) (Shao et al.
-2024), a state-of-the-art reinforcement learning algorithm for fine-tuning LLMs.
-  • We evaluate our GRPO-based method on mathematical problem solving, poetry generation, and the tasks included in NoveltyBench (Zhang et al.
-2025), demonstrating that our scoring metric correlates with domain-specific measures of value and originality.
-Moreover, we show that our approach enhances both the quality and diversity of generated outputs, positioning it as a strong candidate for creativity-focused applications of current foundation models.
-Related Work
-Information Theory and Creativity
+  • We discuss how the score can be practically computed in the case of autoregressive models, and how it can be used as a reward with Group Relative Policy Optimization (GRPO) (Shao et al., 2024), a state-of-the-art reinforcement learning algorithm for fine-tuning LLMs.
+  • We evaluate our GRPO-based method on mathematical problem solving, poetry generation, and the tasks included in NoveltyBench (Zhang et al., 2025), demonstrating that our approach can enhance both the quality and diversity of generated outputs, positioning it as a strong candidate for creativity-focused applications of current foundation models.
+2 Related Work
+2.1 Information Theory and Creativity
 The quest to provide a mathematical and computational definition of creativity has been a significant focus in recent decades.
-Numerous methods have been developed to define various dimensions or attributes for evaluating the creativity of AI-generated products (see, for example, Franceschelli and Musolesi 2024a).
+Numerous methods have been developed to define various dimensions or attributes for evaluating the creativity of AI-generated products (see, for example, Franceschelli & Musolesi, 2024a).
 However, these methods are often domain-specific and typically require substantial human effort to implement and assess.
-In contrast, solutions based on information theory (Shannon 1948; Cover 1999) offer a more universally applicable approach.
+In contrast, solutions based on information theory (Shannon, 1948; Cover, 1999) offer a more universally applicable approach.
 Information-theoretic methods can quantify creativity by measuring the novelty and complexity of generated outputs, without the need for extensive human intervention, making them suitable for a wide range of domains.
-Bayesian surprise (Baldi and Itti 2010), i.e., the divergence between a prior and a posterior belief, has been extensively used to measure either novelty (França et al.
-2016; Varshney et al.
-2019) or surprise (Mazzaglia et al.
-2022; Schmidhuber 2010).
+Bayesian surprise (Baldi & Itti, 2010), i.e., the divergence between a prior and a posterior belief, has been extensively used to measure different shades of originality, such as novelty (França et al., 2016; Varshney et al., 2019) and surprise (Mazzaglia et al., 2022; Schmidhuber, 2010).
 Nevertheless, Varshney (2019) demonstrated that there is a mathematical limit for Bayesian surprise when combined with quality measures.
-Surprisal (Tribus 1961), i.e., Shannon’s self-information, has also been used (Bunescu and Uduehi 2019; Fernandez Monsalve, Frank, and Vigliocco 2012); Barto, Mirolli, and Baldassarre (2013) extensively discuss surprisal and Bayesian surprise, and how novelty differs from them.
-Crucially, in the context of RL, surprisal has been used as a form of intrinsic motivation to encourage the agent to explore more (Achiam and Sastry 2016).
+Surprisal (Tribus, 1961), i.e., Shannon’s self-information, has also been used (Bunescu & Uduehi, 2019; Fernandez Monsalve et al., 2012); Barto et al.
+(2013) extensively discuss surprisal, Bayesian surprise, and novelty.
+Crucially, in the context of RL, surprisal has been used as a form of intrinsic motivation to encourage the agent to explore more (Achiam & Sastry, 2016).
 Sun et al.
-(2025) apply this idea to improve exploration in RLHF (Christiano et al.
-2017).
+(2025) apply this idea to improve exploration in RLHF (Christiano et al., 2017).
+A similar strategy can be applied during LLM fine-tuning, either by explicitly maximizing the model’s perplexity (Dai et al., 2025) or by decoupling entropy and cross-entropy from KL regularization and assigning greater weight to the former (Slocum et al., 2025).
 Burns (2006) proposes to use entropy for expectation and violation, plus posterior probability for explanation in the context of aesthetic experience.
-Additionally, mutual information has been applied to neural conversation models to improve both diversity and appropriateness (Li et al.
-2016).
+Additionally, mutual information has been applied to neural conversation models to improve both diversity and appropriateness (Li et al., 2016).
 However, all these existing approaches are not able to capture and simultaneously optimize value and originality at the same time.
-LLMs and Creativity
-Since the introduction of GPT models (Brown et al.
-2020; OpenAI 2023) and their competitors (e.g., Touvron et al.
-2023), researchers have been keenly exploring the potential for LLMs to exhibit creativity and the methods to achieve this (Franceschelli and Musolesi 2024b).
-For example, human creativity tests like the Alternate Uses Test have been employed to evaluate the creativity of LLMs (Stevenson et al.
-2022) and to investigate methods for enhancing their performance (Goes et al.
-2023; Summers-Stay, Voss, and Lukin 2023).
-Porter and Machery (2024) report that non-expert poetry readers already favor AI-generated poems over human-authored ones.
+2.2 LLMs and Creativity
+Since the introduction of GPT models (Brown et al., 2020; OpenAI, 2023) and their competitors (e.g., Touvron et al., 2023), researchers have been keenly exploring the potential for LLMs to exhibit creativity and the methods to achieve this (Franceschelli & Musolesi, 2025).
+For example, human creativity tests like the Alternate Uses Test have been employed to evaluate the creativity of LLMs (Stevenson et al., 2022) and to investigate methods for enhancing their performance (Goes et al., 2023; Summers-Stay et al., 2023).
+Porter & Machery (2024) report that non-expert poetry readers already favor AI-generated poems over human-authored ones.
 In contrast, Davis (2024) argues that ChatGPT’s poetry is incompetent and banal.
-Either way, instead of being used off-the-shelf, LLMs can be fine-tuned to produce more rhyming poems (Popescu-Belis et al.
-2023) or utilized in zero-shot settings to emulate the writing styles of famous authors (Sawicki et al.
-2023).
-It has also been shown that these models can be fine-tuned via RLHF (Christiano et al.
-2017) to write short poems that human evaluators find more creative (Pardinas et al.
-2023).
-Reinforcement learning (RL) can also encourage large language models (LLMs) to produce more diverse outputs (Chung et al.
-2025), and can be leveraged to optimize metrics commonly linked to creative expression (Ismayilzada et al.
-2025).
-Finally, it is possible to leverage quality-diversity algorithms to generate more creative products; these methods can be based on human (Ding et al.
-2023) or AI (Bradley et al.
-2024) feedback to measure the quality of the generated outputs.
-Preliminaries
-Language Modeling
-A $\boldsymbol{\theta}$-parameterized autoregressive language model is a probability distribution $p_{\boldsymbol{\theta}}(\mathbf{x})$ over a variable-length text sequence $\mathbf{x}=(x_{1}\ldots x_{T})$, where $T$ is the sequence length and each token $x_{t}$ is in a finite vocabulary $\mathcal{V}$ of size $N$.
-The probability distribution is factorized as $p_{\boldsymbol{\theta}}(\mathbf{x})=\prod_{t=1}^{T}p_{\boldsymbol{\theta}}(x_{t}|\mathbf{x_{<t}})$, where $\mathbf{x_{<t}}=x_{1}\ldots x_{t-1}$.
+Either way, instead of being used off-the-shelf, LLMs can be fine-tuned to produce more rhyming poems (Popescu-Belis et al., 2023) or utilized in zero-shot settings to emulate the writing styles of famous authors (Sawicki et al., 2023).
+It has also been shown that these models can be fine-tuned via RLHF (Christiano et al., 2017) to write short poems that human evaluators find more creative (Pardinas et al., 2023).
+Reinforcement learning (RL) can also encourage large language models (LLMs) to produce more diverse outputs (Chung et al., 2025), and can be leveraged to optimize metrics commonly linked to creative expression (Ismayilzada et al., 2025).
+Finally, it is possible to leverage quality-diversity algorithms to generate more creative products; these methods can be based on human (Ding et al., 2023) or AI (Bradley et al., 2024) feedback to measure the quality of the generated outputs.
+3 Preliminaries
+3.1 Language Modeling
+A $\bm{\theta}$-parameterized autoregressive language model is a probability distribution $p_{\bm{\theta}}(\mathbf{x})$ over a variable-length text sequence $\mathbf{x}=(x_{1}\ldots x_{T})$, where $T$ is the sequence length and each token $x_{t}$ is in a finite vocabulary $\mathcal{V}$ of size $N$.
+The probability distribution is factorized as $p_{\bm{\theta}}(\mathbf{x})=\prod_{t=1}^{T}p_{\bm{\theta}}(x_{t}|\mathbf{x_{<t}})$, where $\mathbf{x_{<t}}=x_{1}\ldots x_{t-1}$.
 The language model is usually trained to maximize the likelihood of the true distribution $p^{*}(\mathbf{x})$ for any $\mathbf{x}$ from a reference dataset (the training set).
@@
 While this makes such a model immediately capable of scoring the probability of a given text, it also allows for the generation of new sentences.
-Given a conditional input (prompt) $\mathbf{z}=(z_{1}\ldots z_{L})$, we can decode $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})$ as the continuation of $\mathbf{z}$, i.e., through the factorized representation $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})=\prod_{t=1}^{T}p_{\boldsymbol{\theta}}(x_{t}|\mathbf{x_{<t}},\mathbf{z})$.
-Reinforcement Learning for Language Models
-Due to its adherence to the formal framework of Markov decision processes (Sutton and Barto 2018), RL can be used as a solution to the generative modeling problem in the case of autoregressive tasks such as text generation (Bachman and Precup 2015).
+Given a conditional input (prompt) $\mathbf{z}=(z_{1}\ldots z_{L})$, we can decode $p_{\bm{\theta}}(\mathbf{x}|\mathbf{z})$ as the continuation of $\mathbf{z}$, i.e., through the factorized representation $p_{\bm{\theta}}(\mathbf{x}|\mathbf{z})=\prod_{t=1}^{T}p_{\bm{\theta}}(x_{t}|\mathbf{x_{<t}},\mathbf{z})$.
+3.2 Reinforcement Learning for Language Models
+Due to its adherence to the formal framework of Markov decision processes (Sutton & Barto, 2018), RL can be used as a solution to the generative modeling problem in the case of autoregressive tasks such as text generation (Bachman & Precup, 2015).
 The LLM plays the role of the agent, and each generated token represents an action $a_{t}$.
@@
 Within this framework, any policy-based method can be employed to train or fine-tune the LLM to optimize a given objective.
-Indeed, RL facilitates the use of non-differentiable reward functions, enabling the optimization of test-time metrics, domain-specific targets, and human preferences (Franceschelli and Musolesi 2024c).
-A widely used RL algorithm for fine-tuning LLMs is Proximal Policy Optimization (PPO) (Schulman et al. 2017), which aims to maximize the following objective:
-$\mathcal{J}(\boldsymbol{\theta})=\mathbb{E}_{t}\!\left[\min(r_{t}(\boldsymbol{\theta})\hat{A}_{t},\mathtt{clip}(r_{t}(\boldsymbol{\theta}),1-\epsilon,1+\epsilon)\hat{A}_{t})\right]$ (1)
-where $r_{t}(\boldsymbol{\theta})=\frac{\pi_{\boldsymbol{\theta}}(\mathbf{x}_{t}|\mathbf{z},\mathbf{x}_{<t})}{\pi_{\boldsymbol{\theta}_{old}}(\mathbf{x}_{t}|\mathbf{z},\mathbf{x}_{<t})}$ with $\pi_{\boldsymbol{\theta}}$ and $\pi_{\boldsymbol{\theta}_{old}}$ denoting the current and old policy models, repectively; $\mathbf{x}$ is the output sampled from the old policy given the prompt $\mathbf{z}$; and $\epsilon$ is a clipping factor used to stabilize training.
-The advantage $\hat{A}_{t}$ is usually computed through Generalized Advantage Estimation (Schulman et al.
-2016) based on the full rewards $R(\mathbf{z},\mathbf{x})=r(\mathbf{z},\mathbf{x})-\beta\log\frac{\pi_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})}{\pi_{ref}(\mathbf{x}|\mathbf{z})}$, thus integrating a KL penalty with respect to a reference model (usually, the same model before fine-tuning), and a learned value function $v_{\boldsymbol{\phi}}(\mathbf{s})$.
+Indeed, RL facilitates the use of non-differentiable reward functions, enabling the optimization of test-time metrics, domain-specific targets, and human preferences (Franceschelli & Musolesi, 2024b).
+A widely used RL algorithm for fine-tuning LLMs is Proximal Policy Optimization (PPO) (Schulman et al., 2017), which aims to maximize the following objective:
+$\mathcal{J}(\bm{\theta})=\hat{\mathbb{E}}_{t}\!\left[\min(r_{t}(\bm{\theta})\hat{A}_{t},\mathtt{clip}(r_{t}(\bm{\theta}),1-\epsilon,1+\epsilon)\hat{A}_{t})\right]$ (1)
+where $r_{t}(\bm{\theta})=\frac{\pi_{\bm{\theta}}(\mathbf{x}_{t}|\mathbf{z},\mathbf{x}_{<t})}{\pi_{\bm{\theta}_{old}}(\mathbf{x}_{t}|\mathbf{z},\mathbf{x}_{<t})}$ with $\pi_{\bm{\theta}}$ and $\pi_{\bm{\theta}_{old}}$ denoting the current and old policy models, repectively; $\mathbf{x}$ is the output sampled from the old policy given the prompt $\mathbf{z}$; and $\epsilon$ is a clipping factor used to stabilize training.
+The advantage $\hat{A}_{t}$ is usually computed through Generalized Advantage Estimation (Schulman et al., 2016) based on the full rewards $R(\mathbf{z},\mathbf{x})=r(\mathbf{z},\mathbf{x})-\beta\log\frac{\pi_{\bm{\theta}}(\mathbf{x}|\mathbf{z})}{\pi_{ref}(\mathbf{x}|\mathbf{z})}$, thus integrating a KL penalty with respect to a reference model (usually, the same model before fine-tuning), and a learned value function $v_{\bm{\phi}}(\mathbf{s})$.
 However, learning such a value function is computationally intensive, and its training is complicated by the fact that only the last state is scored by the reward function.
 Moreover, the inclusion of the KL penalty as an auxiliary reward term adds complexity to the advantage estimation process.
-To address these issues, Group Relative Policy Optimization (GRPO) (Shao et al.
-2024) has been introduced.
-GRPO directly adds a KL divergence term $-\beta D_{KL}(\pi_{\boldsymbol{\theta}}||\pi_{ref})$ to the loss rather than to the single rewards, and especially obviates the need for a value function approximator by using the average reward of multiple sampled outputs as the baseline:
+To address these issues, Group Relative Policy Optimization (GRPO) (Shao et al., 2024) has been introduced.
+GRPO directly adds a KL divergence term $\beta D_{KL}(\pi_{\bm{\theta}}||\pi_{ref})$ to the loss rather than to the single rewards, and especially obviates the need for a value function approximator by using the average reward of multiple sampled outputs as the baseline:
 $\hat{A}_{i,t}=\frac{r_{i}-\mathtt{mean}(\mathbf{r})}{\mathtt{std}(\mathbf{r})},$ (2)
 with $\mathbf{r}=\{r_{1},r_{2},...,r_{G}\}$ as the list of rewards received by each of the $G$ outputs sampled from the same prompt.
-A Context-Based Score for Valuable and Original Generation
+4 A Context-Based Score for Valuable and Original Generation
 Our goal is to derive a score that is able to quantify both value and originality at the same time.
@@
 More specifically, we start from the (point-wise) mutual information between two variables $x$ and $y$:
-$I(x,y)=h(x)-h(x|y)=h(y)-h(y|x)$ (3)
-where the entropy is $h(a)=-\log p(a)$, therefore:
-$I(x,y)$ $=\log p(x|y)-\log p(x)$ (4)
-$=\log p(y|x)-\log p(y).$
+I(x,y)=h(x)h(x|y)=h(y)h(y|x)I(x,y)=h(x)-h(x|y)=h(y)-h(y|x) (3)
+where the self-information is $h(a)=-\log p(a)$, therefore:
+$I(x,y)=\log p(x|y)-\log p(x)=\log p(y|x)-\log p(y).$ (4)
 Let us now assume $x$ to be our input vector $\mathbf{x}$ and $y$ our output vector $\mathbf{y}$, obtaining:
@@
 By applying the Bayes theorem, i.e., $\log p(a|b)=\log p(b|a)+\log p(a)-\log p(b)$, we can substitute the $\log p(\mathbf{y})$ term as follows:
-$I(\mathbf{x},\mathbf{y},\lambda_{1},\lambda_{2})$ $=$ $\lambda_{1}\log p(\mathbf{y}|\mathbf{x})-\lambda_{2}\log p(\mathbf{y}|\mathbf{x})-$ (7)
-$\lambda_{2}\log p(\mathbf{x})+\lambda_{2}\log p(\mathbf{x}|\mathbf{y})$
-$=$ $(\lambda_{1}-\lambda_{2})\log p(\mathbf{y}|\mathbf{x})+$
-$\lambda_{2}\log p(\mathbf{x}|\mathbf{y})-\lambda_{2}\log p(\mathbf{x}).$
+$I(\mathbf{x},\mathbf{y},\lambda_{1},\lambda_{2})$ $=$ $\lambda_{1}\log p(\mathbf{y}|\mathbf{x})-\lambda_{2}\log p(\mathbf{y}|\mathbf{x})-\lambda_{2}\log p(\mathbf{x})+\lambda_{2}\log p(\mathbf{x}|\mathbf{y})$ (7)
+$=$ $(\lambda_{1}-\lambda_{2})\log p(\mathbf{y}|\mathbf{x})+\lambda_{2}\log p(\mathbf{x}|\mathbf{y})-\lambda_{2}\log p(\mathbf{x}).$
 Since our goal is to find the optimal $\mathbf{y}$ for a given $\mathbf{x}$, the last term can be ignored.
@@
 In other words, the optimal $\mathbf{y}^{*}$ must be unexpected and diverse from $p(\mathbf{y}|\mathbf{x})$, but it must also be explainable by $\mathbf{x}$.
-While $-\log p(\mathbf{y}|\mathbf{x})$, commonly known as surprisal (Tribus 1961), is widely used to measure diversity and surprise (Barto, Mirolli, and Baldassarre 2013), the other term, $\log p(\mathbf{x}|\mathbf{y})$, can be used to measure value and effectiveness.
-Indeed, if the request (e.g., a problem or a task) can be easily predicted from the outcome, the outcome must be a (good) example of that task or a correct solution for that problem.
-While other prominent definitions, such as Boden’s, seek to include a third requirement for creativity, i.e., novelty (Boden 2003), we believe that by considering only a specific context, novelty and surprise become indistinguishable.
-In particular, being novel in a specific context $\mathbf{x}$ means doing something different from what has been previously experienced under $\mathbf{x}$ by the creator.
-However, since a self-supervised-trained model can be seen as a compression of training data (Franceschelli, Cevenini, and Musolesi 2024), $-\log p(\mathbf{y}|\mathbf{x})$ becomes not only a measure of unexpectedness but also of novelty.
-Therefore, we choose to refer to this term as originality, as it encompasses both novelty and surprise.
-On the other hand, $\log p(\mathbf{x}|\mathbf{y})$ is less ambiguous and maps directly into value.
+$\log p(\mathbf{y}|\mathbf{x})$, commonly known as surprisal (Tribus, 1961), is widely used to measure diversity and surprise (Barto et al., 2013), and adheres to the first requirement from the standard definition of creativity by Runco & Jaeger (2012), i.e., originality.
+Conversely, $\log p(\mathbf{x}|\mathbf{y})$ can be used to measure value or effectiveness, the second requirement of the definition.
+If the request (e.g., a problem or task) can be inferred from the outcome, the latter constitutes an appropriate instance of that task or a correct, useful solution to that problem (e.g., if the request is for a sonnet or a sci-fi screenplay, the generated artifact is valuable if identified as a poem satisfying the metrical constraints of a sonnet or as a screenplay adhering to a sci-fi theme).
 In summary, the CoVO (Context-based Value and Originality) score for a target $\mathbf{y}$ given a source $\mathbf{x}$ on a reference probability distribution $p$ is formally defined as:
 $s_{CoVO}(\mathbf{x},\mathbf{y},p)=\underbrace{\lambda_{v}\log p(\mathbf{x}|\mathbf{y})}_{\textrm{Value}}\underbrace{-\lambda_{o}\log p(\mathbf{y}|\mathbf{x})}_{\textrm{Originality}}$ (9)
-Implementation and Optimization with Autoregressive Models
+5 Implementation and Optimization with Autoregressive Models
 We now discuss the implementation of the CoVO score with autoregressive models.
-Using the notation introduced above, in the context of a $\boldsymbol{\theta}$-parameterized LLM, $p(\mathbf{y}|\mathbf{x})$ can be expressed as $\prod_{t=1}^{T}p_{\boldsymbol{\theta}}(y_{t}|\mathbf{y}_{<t},\mathbf{x})$.
+Using the notation introduced above, in the context of a $\bm{\theta}$-parameterized LLM, $p(\mathbf{y}|\mathbf{x})$ can be expressed as $\prod_{t=1}^{T}p_{\bm{\theta}}(y_{t}|\mathbf{y}_{<t},\mathbf{x})$.
 However, considering just the product of all the conditioned probabilities for an optimization problem would lead to preferring shorter sequences.
-To avoid this, we propose to use the $T$-th root: $T]{\prod_{t=1}^{T}p_{\boldsymbol{\theta}}(y_{t}|\mathbf{y}_{<t},\mathbf{x})}$.
+To avoid this, we propose to use the $T$-th root: $\sqrt[T]{\prod_{t=1}^{T}p_{\bm{\theta}}(y_{t}|\mathbf{y}_{<t},\mathbf{x})}$.
 By leveraging the properties of the logarithm, we obtain:
-$s_{CoVO}^{AR}$ $=\lambda_{v}\frac{\sum_{i=1}^{|\mathbf{x}|}\log p_{\boldsymbol{\theta}}(x_{i}|\mathbf{x}_{<i},\mathbf{y})}{|\mathbf{x}|}$ (10)
-$-\lambda_{o}\frac{\sum_{j=1}^{|\mathbf{y}|}\log p_{\boldsymbol{\theta}}(y_{j}|\mathbf{y}_{<j},\mathbf{x})}{|\mathbf{y}|}.$
-It is worth noting that the vocabulary of an LLM can be extremely large, which can cause $p_{\boldsymbol{\theta}}(a|b)$ to be small even when $a$ is the most probable event given $b$.
-In particular, when an LLM generates $\mathbf{y}$ given $\mathbf{x}$ and then evaluates both $p_{\boldsymbol{\theta}}(\mathbf{y}|\mathbf{x})$ and $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{y})$, this can lead to a significant discrepancy between the magnitude of value and diversity.
-Since $\mathbf{y}$ has been sampled from a model based on $p_{\boldsymbol{\theta}}$, its probability would be high by definition.
+$s_{CoVO}^{AR}=\lambda_{v}\frac{\sum_{i=1}^{|\mathbf{x}|}\log p_{\bm{\theta}}(x_{i}|\mathbf{x}_{<i},\mathbf{y})}{|\mathbf{x}|}-\lambda_{o}\frac{\sum_{j=1}^{|\mathbf{y}|}\log p_{\bm{\theta}}(y_{j}|\mathbf{y}_{<j},\mathbf{x})}{|\mathbf{y}|}.$ (10)
+It is worth noting that the vocabulary of an LLM can be extremely large, which can cause $p_{\bm{\theta}}(a|b)$ to be small even when $a$ is the most probable event given $b$.
+In particular, when an LLM generates $\mathbf{y}$ given $\mathbf{x}$ and then evaluates both $p_{\bm{\theta}}(\mathbf{y}|\mathbf{x})$ and $p_{\bm{\theta}}(\mathbf{x}|\mathbf{y})$, this can lead to a significant discrepancy between the magnitude of value and diversity.
+Since $\mathbf{y}$ has been sampled from a model based on $p_{\bm{\theta}}$, its probability would be high by definition.
 However, there may be various ways (possibly through synonyms) to define $\mathbf{y}$, leading to a smaller probability of $\mathbf{x}$.
-Inspired by Macedo, Reisenzein, and Cardoso (2004), we propose to counteract this problem by normalizing $p_{\boldsymbol{\theta}}(a|b)$ via $n^{\prime}=\frac{n-n_{min}}{n_{max}-n_{min}}$.
-For probabilities, $n_{min}=0$, while $n_{max}=\max p_{\boldsymbol{\theta}}(b)$, thus obtaining the overall mapping for $p_{\boldsymbol{\theta}}$: $\frac{p_{\boldsymbol{\theta}}(y_{t}|\mathbf{y}_{<t},\mathbf{x})}{\max p_{\boldsymbol{\theta}}(\mathbf{y}_{<t},\mathbf{x})}$.
+Inspired by Macedo et al.
+(2004), we propose to counteract this problem by normalizing $p_{\bm{\theta}}(a|b)$ via $n^{\prime}=\frac{n-n_{min}}{n_{max}-n_{min}}$.
+For probabilities, $n_{min}=0$, while $n_{max}=\max_{v\in\mathcal{V}}p_{\bm{\theta}}(b)$, thus obtaining the overall mapping for $p_{\bm{\theta}}$: $\frac{p_{\bm{\theta}}(y_{t}|\mathbf{y}_{<t},\mathbf{x})}{\max_{v\in\mathcal{V}}p_{\bm{\theta}}(\mathbf{y}_{<t},\mathbf{x})}$.
 Once again, by applying the properties of logarithms, we obtain:
-$s_{CoVO}^{AR_{norm}}(\mathbf{x},\mathbf{y},p_{\boldsymbol{\theta}})=\lambda_{v}s_{v}(\mathbf{x},\mathbf{y},p_{\boldsymbol{\theta}})+\lambda_{o}s_{o}(\mathbf{x},\mathbf{y},p_{\boldsymbol{\theta}})=$ (11)
-$\lambda_{v}\frac{\sum_{i=1}^{|\mathbf{x}|}(\log p_{\boldsymbol{\theta}}(x_{i}|\mathbf{x}_{<i},\mathbf{y})-\max\log p_{\boldsymbol{\theta}}(\mathbf{x}_{<i},\mathbf{y}))}{|\mathbf{x}|}-$
-$\lambda_{o}\frac{\sum_{j=1}^{|\mathbf{y}|}(\log p_{\boldsymbol{\theta}}(y_{j}|\mathbf{y}_{<j},\mathbf{x})-\max\log p_{\boldsymbol{\theta}}(\mathbf{y}_{<j},\mathbf{x}))}{|\mathbf{y}|}.$
-Calculating $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{y})$ is not trivial.
+$s_{CoVO}^{AR_{norm}}(\mathbf{x},\mathbf{y},p_{\bm{\theta}})=\lambda_{v}s_{v}(\mathbf{x},\mathbf{y},p_{\bm{\theta}})+\lambda_{o}s_{o}(\mathbf{x},\mathbf{y},p_{\bm{\theta}})=$ (11)
+$\lambda_{v}\frac{\sum_{i=1}^{|\mathbf{x}|}(\log p_{\bm{\theta}}(x_{i}|\mathbf{x}_{<i},\mathbf{y})-\max_{v\in\mathcal{V}}\log p_{\bm{\theta}}(\mathbf{x}_{<i},\mathbf{y}))}{|\mathbf{x}|}$-
+$\lambda_{o}\frac{\sum_{j=1}^{|\mathbf{y}|}(\log p_{\bm{\theta}}(y_{j}|\mathbf{y}_{<j},\mathbf{x})-\max_{v\in\mathcal{V}}\log p_{\bm{\theta}}(\mathbf{y}_{<j},\mathbf{x}))}{|\mathbf{y}|}.$
+Calculating $p_{\bm{\theta}}(\mathbf{x}|\mathbf{y})$ is not trivial.
 Since LLMs are trained to complete text sequences, it is unlikely that they would generate the source text immediately after the target text (which, we should remember, is generated immediately after the source text).
-To address this, we consider an approximation $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{y}^{\prime})$, where $\mathbf{y}^{\prime}=\mathbf{y}+\mathbf{q}$.
+To address this, we consider an approximation $p_{\bm{\theta}}(\mathbf{x}|\mathbf{y}^{\prime})$, where $\mathbf{y}^{\prime}=\mathbf{y}+\mathbf{q}$.
 Here, $\mathbf{q}$ represents an additional question, such as “How would you describe this text?” or a similar formulation designed solely to increase the likelihood of generating the source text $\mathbf{x}$ (as well as alternative sources).
@@
 Then, the model can be trained with any policy gradient method.
-Our experiments leverage GRPO (Shao et al.
-2024), which is a state-of-the-art choice for training language models.
+Our experiments leverage GRPO (Shao et al., 2024), which is a state-of-the-art choice for training language models.
 As introduced above, GRPO adds a per-token KL divergence term to the loss rather than to the single rewards.
-Usually, the KL divergence is approximated with the following estimator (Schulman 2020): $r_{ref}(\boldsymbol{\theta})-\log r_{ref}(\boldsymbol{\theta})-1$, where $r_{ref}(\boldsymbol{\theta})=\frac{\pi_{ref}}{\pi_{\boldsymbol{\theta}}}$ and $\pi_{ref}$ is the reference policy, i.e., the model before GRPO training.
-In particular, GRPO aims to minimize the KL divergence, thus the second term $-\log r_{ref}=-\log\pi_{ref}+\log\pi_{\boldsymbol{\theta}}$ can be seen as made of two optimization problems: the minimization of the originality component $-\log\pi_{ref}$ from Equation 9, and the maximization of surprisal, or self-information, $-\log\pi_{\boldsymbol{\theta}}$ under the current model.
+Usually, the KL divergence is approximated with the following estimator (Schulman, 2020): $r_{ref}(\bm{\theta})-\log r_{ref}(\bm{\theta})-1$, where $r_{ref}(\bm{\theta})=\frac{\pi_{ref}}{\pi_{\bm{\theta}}}$ and $\pi_{ref}$ is the reference policy, i.e., the model before GRPO training.
+In particular, GRPO aims to minimize the KL divergence, thus the second term $\log r_{ref}=-\log\pi_{ref}+\log\pi_{\bm{\theta}}$ can be seen as made of two optimization problems: the minimization of the originality component $\log\pi_{ref}$ from Equation 9, and the maximization of surprisal, or self-information, $\log\pi_{\bm{\theta}}$ under the current model.
 In other words, this second term somehow trades off a portion of the originality component under the reference model (proportional to the $\beta\ll 1$ coefficient) with the surprisal under the current model.
 However, the first term of the KL approximation keeps the two policies closer, preventing the trained one from deviating too much and potentially disrupting the impact of the originality component in favor of the value component.
-Experiments
+6 Experiments
 We evaluate the effectiveness of our RL strategy through three case studies: poetry generation, mathematical problem resolution, and the tasks included in NoveltyBench [^1] The code and results of the experiments can be found at
@@
 Both methods assume $\lambda_{v}=\lambda_{o}=1.0$.
-Poetry Generation
+While it is common to induce diversity at the sampling level (e.g., through min-$p$ (Minh et al., 2025) or conformative decoding (Peeperkorn et al., 2025)), we restrict our baselines to the original model and, if available, a model tuned solely on the environmental reward.
+Our approach is orthogonal to the chosen sampling strategy and can benefit from more advanced methods.
+In our evaluation, we aim at verifying whether our reward scheme can increase value and originality, regardless of how the output is sampled.
+6.1 Poetry Generation
 Method In-distribution Out-of-distribution
 Corr. $\uparrow$ Metric (L/S) $\uparrow$ T-LCS $\downarrow$ Tone $\uparrow$ Corr. $\uparrow$ Metric (L/S) $\uparrow$ T-LCS $\downarrow$ Tone $\uparrow$
-Meta-Llama-3-8B 0.9870.9870.987 0.3000.3000.300 / 0.1770.1770.177 6.0676.0676.067 / 191919 $0.664_{\pm 0.069}$ 1.0001.0001.000 0.4440.4440.444 / 0.1320.1320.132 5.8535.8535.853 / 686868 $0.621_{\pm 0.063}$
-+ CoVO 1.0001.0001.000 0.2000.2000.200 / 0.0840.0840.084 4.8804.8804.880 / 0777 $0.749_{\pm 0.055}$ 0.9600.9600.960 0.2670.2670.267 / 0.0460.0460.046 4.9334.9334.933 / 0666 $0.661_{\pm 0.060}$
-+ CoVO + KL 0.9870.9870.987 0.3330.3330.333 / 0.0660.0660.066 5.1075.1075.107 / 0999 $0.745_{\pm 0.057}$ 0.9730.9730.973 0.2890.2890.289 / 0.1170.1170.117 5.1605.1605.160 / 222222 $0.682_{\pm 0.059}$
-Table 1: Aggregate results of CoVO scores at inference time considering both training prompts (left) and testing prompts (right).
+Meta-Llama-3-8B $0.987$ $0.300$ / $0.177$ $6.067$ / 1919 $0.664_{\pm 0.069}$ $1.000$ $0.444$ / $0.132$ $5.853$ / $68$ $0.621_{\pm 0.063}$
++ CoVO $1.000$ $0.200$ / $0.084$ $4.880$ / 077 $0.749_{\pm 0.055}$ $0.960$ $0.267$ / $0.046$ $4.933$ / 066 $0.661_{\pm 0.060}$
++ CoVO + KL $0.987$ $0.333$ / $0.066$ $5.107$ / 099 $0.745_{\pm 0.057}$ $0.973$ $0.289$ / $0.117$ $5.160$ / $22$ $0.682_{\pm 0.059}$
+Table 1: Aggregate results of generated poems considering both training prompts (left) and testing prompts (right).
 Scores on the poetical metrics are reported at the line level (L) and syllable level (S) and only consider requests for styles with specific metrical properties.
@@
 The mean and the 95% confidence interval are reported for tone adherence.
+Figure 2: The distribution of value and originality (according to our scores) for the in-distribution and out-of-distribution poems generated by the baseline and our two methods.
 Experimental Setup.
-[…]
+The first set of experiments concerns a very common creative task, aiming to teach the LLM to generate poems that are both more original and valuable.
+More specifically, we follow the approach outlined by Bradley et al.
+(2024) and instruct the model to write a poem in a particular style and tone.
+We consider the Meta-Llama-3-8B model (Grattafiori et al., 2024) as our pre-trained agent.
+Since we do not use the instruction-tuned model, we prompt it with some few-shot examples of the task to make it more likely to produce the desired output in the desired form (the full prompt is reported in Appendix A, together with the full training parameters).
+Instead of fine-tuning the entire network, we consider Low-Rank Adaptation (LoRA) (Hu et al., 2022).
+The original model is also used to compute the score.
+All sampling happens with top-$k$ ($k=50$) and at a temperature of $1.0$.
+We perform a quantitative evaluation where we compute poetical metrics for quality (lexical correctness of poems, adherence to line- and syllable-level constraints, and tone adherence to the request through zero-shot classification (Yin et al., 2019) with bart-large-mnli model (Lewis et al., 2020)) and for originality (accidental reproduction of existing poems).
+For the latter, we define a Token-based Longest Common Substring (T-LCS) score, and we use it by comparing generated poems with a reference dataset of approx.
+84k public-domain poems extracted from Project Gutenberg (please refer to Appendix B for a first presentation of our GutenVerse dataset).
+While a generated poem can be an accidental reproduction of a protected work or a different kind of text (e.g., a song), we believe it can provide a useful evaluation tool to understand the general degree of originality.
 Experimental Results.
@@
 Figure 2 reports the value and originality according to Equation 11 under the pre-trained model.
-While the two methods do not significantly differ from the baseline (which is possibly due to the opposite forces of value and originality (Varshney 2019)), we again see that the presence of KL leads to slightly higher value, while its absence leads to slightly higher originality.
+While the two methods do not significantly differ from the baseline (which is possibly due to the opposite forces of value and originality (Varshney, 2019)), we again see that the presence of KL leads to slightly higher value, while its absence leads to slightly higher originality.
 However, aggregated scores, such as those presented here, might be insufficient.
-For a more complete overview, we also conducted a fine-grained analysis of the generated poems in the Technical Appendix.
-Figure 2: The distribution of value and originality (according to our scores) for the in-distribution and out-of-distribution poems generated by the baseline and our four methods.
-Math Problem Resolution
+For a more complete overview, we also conducted a fine-grained analysis of the generated poems in Appendix C.
+6.2 Math Problem Resolution
 Experimental Setup.
 The second set of experiments concerns a more practical and quantitative task, as it aims to teach the LLM to solve mathematical problems through more diverse procedures.
-In particular, we focus on the Mistral-based (Jiang et al.
-2023) MetaMath-Mistral-7B model, i.e., fine-tuned with self-supervised learning on the MetaMathQA (Yu et al.
-2024).
+In particular, we focus on the Mistral-based (Jiang et al., 2023) MetaMath-Mistral-7B model, i.e., fine-tuned with self-supervised learning on the MetaMathQA (Yu et al., 2024).
 It is a dataset of textual math questions paired with responses where the numerical answer is easily separable from the textual procedure.
-While the entire set contains 395k entries, making an additional training epoch too expensive, MetaMathQA is composed of entries from two different training sets, then augmented with various techniques: GSM8K (Cobbe et al.
-2021) and MATH (Hendrycks et al.
-2021).
+While the entire set contains 395k entries, making an additional training epoch too expensive, MetaMathQA is composed of entries from two different training sets, then augmented with various techniques: GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021).
 Since we are only interested in the questions, we limit our training to a single epoch over those datasets.
@@
 We separate the procedure and the answer from each solution to train our model and use the numerical answer to check the correctness of the predicted solution.
-Because of this, the RL problem can be formulated considering up to two rewards: our CoVO score computed on the procedure and an extrinsic reward based on the correctness of the answer.
-As for the previous case study, instead of fine-tuning the entire model, we adopt a more parameter-efficient strategy with LoRA and use the original model to perform the CoVO score computation.
-Following (Yu et al.
-2024), the outputs are obtained with a greedy strategy.
+The RL problem can then be formulated considering up to two rewards: our CoVO score computed on the procedure and a verifiable, extrinsic reward based on the correctness of the answer.
+Instead of fine-tuning the entire model, we adopt a more parameter-efficient strategy with LoRA and use the original model to perform the CoVO score computation.
+Following Yu et al.
+(2024), the outputs are obtained with a greedy strategy.
 The evaluation considers both GSM8K and MATH test sets (limited to the entries with a tokenized length of question and answer smaller than 512, i.e., all 1319 entries for GSM8K and 4546 out of 5000 for MATH).
-We compute the percentage of correct solutions together with two diversity metrics: expectation-adjusted distinct N-grams (EAD) (Liu et al.
-2022) and sentence embedding cosine similarity (SBERT) (Hong et al.
-2024), which should measure syntactical and semantical diversity, respectively (Kirk et al.
-2024).
+We compute the percentage of correct solutions together with two diversity metrics: expectation-adjusted distinct N-grams (EAD) (Liu et al., 2022) and sentence embedding cosine similarity (SBERT) (Hong et al., 2024), which should measure syntactical and semantical diversity, respectively (Kirk et al., 2024).
 EAD counts the number of distinct N-grams (averaging over $N=1\ldots 5$) across all generated responses and removes the bias toward shorter inputs by scaling the number of distinct tokens based on their expectations.
 The SBERT metric computes the average of the cosine similarity between the embeddings of any possible pairs of outputs and returns 1 minus the similarity.
-This was originally based on Sentence-BERT (Reimers and Gurevych 2019), we employ instead the more recent all-mpnet-base-v2, as suggested by their developers [^2] https://huggingface.co/sentence-transformers/bert-large-nli-stsb-mean-tokens.
+This was originally based on Sentence-BERT (Reimers & Gurevych, 2019), we employ instead the more recent all-mpnet-base-v2, as suggested by their developers [^2] https://huggingface.co/sentence-transformers/bert-large-nli-stsb-mean-tokens.
 Following Kirk et al.
@@
 EAD $\uparrow$ SBERT $\uparrow$ EAD $\uparrow$ SBERT $\uparrow$
-[…]
-[…]
-[…]
-[…]
-[…]
-+ CoVO + Ext $\mathbf{78.33}\%(4)$ 2.03402.03402.0340 $\mathbf{0.6404}$ $0.0628_{\pm.0056}$ $0.0088_{\pm.0011}$
-+ CoVO + Ext + KL $\underline{77.95}\%(4)$ 2.03672.03672.0367 0.64020.64020.6402 $0.0638_{\pm.0057}$ $0.0089_{\pm.0011}$
+MetaMath-Mistral-7B $77.96\%(3)$ $2.0071$ $0.6402$ - -
++ Ext $78.18\%(4)$ $\underline{2.0045}$ $\mathbf{0.6404}$ $\underline{0.0081_{\pm.0021}}$ $\underline{0.0008_{\pm.0002}}$
++ Ext + KL $78.08\%(5)$ $2.0077$ $\underline{0.6401}$ $\underline{0.0096_{\pm.0024}}$ $\underline{0.0011_{\pm.0004}}$
++ CoVO $78.12\%(3)$ $\mathbf{2.0509}$ $0.6403$ $\mathbf{0.0854_{\pm.0062}}$ $\mathbf{0.0118_{\pm.0012}}$
++ CoVO + KL $78.12\%(3)$ $2.0464$ $0.6402$ $\mathbf{0.0879_{\pm.0063}}$ $\mathbf{0.0122_{\pm.0013}}$
++ CoVO + Ext $\mathbf{78.33}\%(4)$ $2.0340$ $\mathbf{0.6404}$ $0628_{\pm.0056} 0.0$$0088_{\pm.0011}$
++ CoVO + Ext + KL $\underline{77.95}\%(4)$ $2.0367$ $0.6402$ $0638_{\pm.0057} 0.0$$0089_{\pm.0011}$
 Table 2: Accuracy and diversity of results for the GSM8k test set.
@@
 EAD $\uparrow$ SBERT $\uparrow$ EAD $\uparrow$ SBERT $\uparrow$
-MetaMath-Mistral-7B 33.55%(483)33.$5$ ( 483 ) 5.72395.72395.7239 0.80320.80320.8032 - -
+MetaMath-Mistral-7B $33.55\%(483)$ $5.7239$ $0.8032$ - -
 + Ext $\underline{33.19}\%(469)$ $\underline{5.7187}$ $\underline{0.8027}$ $\underline{0.0333_{\pm.0029}}$ $\underline{0.0074_{\pm.0008}}$
-[…]
-[…]
-[…]
-+ CoVO + Ext $\mathbf{33.76}\%(503)$ 5.81145.81145.8114 0.80300.80300.8030 $0.1136_{\pm.0047}$ $0.0259_{\pm.0015}$
-[…]
++ Ext + KL $33.56\%(476)$ $5.7517$ $0.8028$ $\underline{0.0345_{\pm.0030}}$ $\underline{0.0075_{\pm.0008}}$
++ CoVO $33.29\%(521)$ $5.8219$ $0.8029$ $\mathbf{0.1457_{\pm.0049}}$ $\mathbf{0.0339_{\pm.0016}}$
++ CoVO + KL $32.79\%(533)$ $\mathbf{5.8442}$ $0.8030$ $\mathbf{0.1479_{\pm.0050}}$ $\mathbf{0.0343_{\pm.0016}}$
++ CoVO + Ext $\mathbf{33.76}\%(503)$ $5.8114$ $0.8030$ $1136_{\pm.0047} 0.0$$0259_{\pm.0015}$
++ CoVO + Ext + KL $33.74\%(497)$ $5.8082$ $0.8031$ $1102_{\pm.0047} 0.0$$0250_{\pm.0015}$
 Table 3: Accuracy and diversity of results for the MATH test set.
@@
 Figure 3: The mean utility (quality) and mean distinct (novelty) scores on NoveltyBench (‘curated’ partition) for the original model and our methods (tuned on the ‘wildchat’ partition).
-NoveltyBench
+6.3 NoveltyBench
 Experimental Setup.
-Finally, we also experiment with NoveltyBench (Zhang et al.
-2025), a very recent benchmark that aims to evaluate the ability of language models to produce multiple distinct and high-quality outputs.
-NoveltyBench contains two sets of prompts thought for eliciting diverse responses: the ‘curated’ partition, with 100 prompts manually curated by the paper’s authors, and the ‘wildchat’ partition, with 1000 prompts sourced from the WildChat-1M dataset (Zhao et al.
-2024).
-The benchmark requires a language model to generate 10 outputs per prompt using a temperature setting of 1.01.01.0, after which it computes scores for novelty and quality.
-To assess novelty, the 10 outputs are grouped into equivalence classes using a fine-tuned DeBERTa model (He et al.
-2021), and the number of distinct classes is reported as the novelty score.
-For the latter, it computes the cumulative utility of the 10 outputs, where the utility is zero if the $i$-th output has the same equivalence class as a precedent output, and a calibrated reward from Skywork-Reward-Gemma-2-27B-v0.2 (Liu et al.
-2024) otherwise (we refer to Zhang et al.
+Finally, we also experiment with NoveltyBench (Zhang et al., 2025), a very recent benchmark that aims to evaluate the ability of language models to produce multiple distinct and high-quality outputs.
+NoveltyBench contains two sets of prompts thought for eliciting diverse responses: the ‘curated’ partition, with 100 prompts manually curated by the paper’s authors, and the ‘wildchat’ partition, with 1000 prompts sourced from the WildChat-1M dataset (Zhao et al., 2024).
+The benchmark requires a language model to generate 10 outputs per prompt using a temperature setting of $1.0$, after which it computes scores for novelty and quality.
+To assess novelty, the 10 outputs are grouped into equivalence classes using a fine-tuned DeBERTa model (He et al., 2021), and the number of distinct classes is reported as the novelty score.
+To assess quality, it computes the cumulative utility of the 10 outputs, where the utility is zero if the $i$-th output has the same equivalence class as a precedent output, and a calibrated reward from Skywork-Reward-Gemma-2-27B-v0.2 (Liu et al., 2024) otherwise (we refer to Zhang et al.
 (2025) for full details).
-To evaluate our approach with NoveltyBench, we fine-tune the Llama-3.2-3B-Instruct model (Grattafiori et al.
-2024) for a single epoch on the ‘wildchat’ partition (the full parameters are reported in the Technical Appendix); then, we compute the novelty and quality scores on the ‘curated’ partition and we compare them with the performance of the original model.
+To evaluate our approach with NoveltyBench, we fine-tune the Llama-3.2-3B-Instruct model (Grattafiori et al., 2024) for a single epoch on the ‘wildchat’ partition (the full parameters are reported in Appendix A); then, we compute the novelty and quality scores on the ‘curated’ partition.
 Experimental Results.
@@
 Moreover, these results underscore the interplay between KL loss and our CoVO score: incorporating the KL penalty tends to improve quality, but at the cost of reduced novelty.
-Discussion
+7 Discussion
 The CoVO score captures properties that are functionally aligned with creativity-relevant aspects of language generation, as demonstrated through both conceptual analysis and empirical results.
@@
 There are a few limitations worth noting.
-First of all, our score represents only a quantifiable approximation of a particular theoretical perspective on creativity, grounded in the dimensions of value and originality.
-It reflects a specific view of the evaluation of creativity based on the generated outputs and does not account for potential alternative theories (for example, arising from different cultures (Lubart 1999)) and perspectives (Rhodes 1961).
-In addition, our experiments are currently limited to only three relatively short-form text generation tasks.
+Firstly, our score represents only a quantifiable approximation of a particular theoretical perspective on creativity, grounded in the dimensions of value and originality.
+For example, value has been considered from the perspective of effectiveness, while other dimensions have been proposed as well (e.g., interestingness (Boden, 1994) or monetary worth (Lepak et al., 2007)); and other classic definitions of creativity, such as that presented in Boden (2003), add a third requirement by splitting originality into novelty and surprise.
+Moreover, our score reflects a specific view of the evaluation of creativity based on the generated outputs and does not account for potential alternative theories (for example, arising from different cultures (Lubart, 1999)) and perspectives (Rhodes, 1961).
+Finally, our experiments are currently limited to only three relatively short-form text generation tasks.
 While their generalizability is supported by the theoretical framework discussed above, the resulting performance was experimentally evaluated for a finite number of scenarios.
-Conclusion
+8 Conclusion
 In this paper, we presented CoVO, a novel score that quantifies the value and originality of neurally-generated text.
-The definition of CoVO is based on analysis of mutual information between the model’s outputs and inputs, and vice versa.
+The definition of CoVO is based on the analysis of mutual information between the model’s outputs and inputs, and vice versa.
 We also proposed an optimization problem where a generative model aims to maximize this score to generate more creative products, and detailed how to use it in language modeling.
 We conducted experiments on poetry generation, math problem solving, and tasks included in NoveltyBench, exploring trade-offs in accuracy vs diversity.
-Effectively balancing value and originality maximization remains an open question, but our score seems to correlate with domain-specific measures appropriately.
+Effectively balancing value and originality maximization remains an open question, but our score seems to relate to domain-specific measures appropriately.
 In addition, fine-tuning to maximize it improves quality- and diversity-related metrics.
-Our research agenda aims to extend our method to other models and tasks, to include inference-level strategies such as creativity-oriented sampling schemes, and to explore its use for evaluation (e.g., in a Best-of-N setting (Stiennon et al.
-2020)) rather than solely for optimization.
+Our research agenda aims to extend our method to other models and tasks, to include inference-level strategies such as creativity-oriented sampling schemes, and to explore its use for evaluation (e.g., in a Best-of-N setting (Stiennon et al., 2020)) rather than solely for optimization.
 We also plan to investigate the definition of additional scores for capturing other potentially relevant aspects of the creative process.
-Despite being costly and inherently constrained (Davis 2024), assessing whether our creativity score aligns with human judgment is another key direction for future work.
+Despite being costly and inherently constrained (Davis, 2024), assessing whether our creativity score aligns with human judgment is another key direction for future work.
 Appendix A Implementation Details
@@
 Parameter Value
+Total batches 100
+Batch size $B$ 4
+Gradient accumulation steps 8
+Max new tokens 256
+Temperature 1.
+Top-$k$ 0
+Optimizer Adam
+Learning rate 1e-5
+Max gradient normalization 100.
+Rank (LoRA) 16
+$\alpha$ parameter (LoRA) 32
+Dropout (LoRA) 0.05
+Training iterations 1
+Scale rewards True
+$\beta$ (when used) 0.05
+Number of generations $G$ 4
+Table 4: Training parameters for poetry generation.
+Parameter Value
 Total epochs 1
@@
 Temperature 1.
-[…]
+Top-$k$ 0
 Optimizer Adam
@@
 Temperature 1.
-[…]
+Top-$k$ 0
 Optimizer Adam
@@
 At inference time we also consider ‘cinquain’, ‘couplet’, ‘free verse’, ‘ode’ or ‘tanka’ as styles and ‘cutting’, ‘nostalgic’, ‘poignant’, ‘solemn’ or ‘whimsical’ as tones.
-Instead, the prompt used for computing $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{y})$ is:
+Instead, the prompt used for computing $p_{\bm{\theta}}(\mathbf{x}|\mathbf{y})$ is:
 Describe the style of the following poem in two words: {prova} I would describe it as a
@@
 Below is an instruction that describes a task. Write a response that appropriately completes the request. ### Instruction: {question} ### Response: Let’s think step by step.
-at inference time. Instead, for computing $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{y})$ we used the following:
+at inference time. Instead, for computing $p_{\bm{\theta}}(\mathbf{x}|\mathbf{y})$ we used the following:
 Below is a response that appropriately completes a request. Write the instruction that describes the task. ### Response: {response} ### Instruction:
@@
 assistant
-where user and assistant are keywords used by the model to identify different roles in the chat. Instead, for computing $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{y})$ we used the following:
+where user and assistant are keywords used by the model to identify different roles in the chat. Instead, for computing $p_{\bm{\theta}}(\mathbf{x}|\mathbf{y})$ we used the following:
 Below is a response that appropriately solves a task. Write the instruction that describes the task. ### Response: {response} ### Instruction:
@@
 To evaluate the accidental reproduction rate of generated poems, we introduce the GutenVerse dataset, which comprises over 84,000 public-domain, English-written poems extracted from Project Gutenberg.
-While generated poems can reproduce different content, e.g., songs or copyrighted material, we believe this can provide a useful indication of how likely a text is original or not.
+While generated poems can reproduce different content, e.g., songs or copyrighted material, we believe this can provide a useful indication of how likely a text is to be original or not.
 To derive our dataset, we started from Gutenberg, dammit [^3] https://github.com/aparrish/gutenberg-dammit/, a corpus of every plaintext file in Project Gutenberg (up until June 2016).
@@
 While it can still contain content that is not poetry (e.g., a table of contents formatted very uncommonly), the poems can be effectively used to measure the overlap between real and generated text.
-We also released a datasheet (Gebru et al. 2018) for the GutenVerse dataset that can be found, together with the code used to create it, at: https://anonymous.4open.science/r/GutenVerse-DD32/.
+We also released a datasheet (Gebru et al., 2018) for the GutenVerse dataset that can be found, together with the code used to create it, at: https://anonymous.4open.science/r/GutenVerse-DD32/.
 Appendix C Detailed Analysis of the Generated Poems